### v2

# A contact topological glossary for non-equilibrium thermodynamics

Abstract
We discuss the occurrence of some notions and results from contact topology in the non-equilibrium thermodynamics. This includes the Reeb chords and the partial order on the space of Legendrian submanifolds.
1 Introduction
The contact geometric formulation of equilibrium thermodynamics is classical and traces back to Gibbs (see [21] for a modern exposition). In that approach, one starts with a manifold $X$ formed by the intensive thermodynamic variables (prescribed external physical parameters). Each macroscopic thermodynamic state is represented by a point in the jet space $J^{1}X=T^{*}X\times{\mathbb{R}}$ that includes the intensive variables, the extensive thermodynamic variables (entropy and the generalized pressures) and the free energy. The starting point of the equilibrium contact thermodynamics is the observation that the set of all thermodynamic equilibria forms a Legendrian submanifold in $J^{1}X$.
In contrast, there exists a variety of contact geometric approaches to mathematical modeling of non-equilibrium thermodynamic processes [4, 11, 12, 14, 15, 16, 17, 18, 19, 20, 23, 28]. In this note, we present yet another approach that examines slow (quasi-static) irreversible processes from the perspective of the partial order on the space of Legendrian submanifolds, a profound geometric structure studied within contact topology [1, 5, 7]; see also [9]. The starting observation is that the first and second laws of thermodynamics impose a constraint: any time-dependent path generated by an out-of-the-equilibrium thermodynamic process must be non-negative with respect to the standard contact form on $J^{1}X$ (see Definition 3.1 below), which is a reformulation of the result in [20]. The main goal of this note is to discuss the consequences of this observation.
We consider three types of non-equilibrium processes that respect the non-negativity of the thermodynamic paths:
(1) In the slow (quasi-static) global [^1] The term “global” here emphasizes that we consider processes defined at all macroscopic equilibrium states of the system, and not only at some of them. processes, the change in the external parameters is so slow that, at each moment in time, the system is in equilibrium. Such processes are given by non-negative families (or paths) of Legendrian equilibrium submanifolds.
(2) In the fast processes, after an abrupt change in the extensive variables, the system converges, as $t\to+\infty$, to the equilibrium corresponding to the new values of the parameters but is not at an equilibrium at the intermediate times $t>0$. Such processes are described by a thermodynamic path driven by a Fokker-Planck equation for the probability density on the microscopic thermodynamic states.
(3) In the ultrafast (instantaneous) processes, the system jumps instantaneously from one equilibrium into another after an abrupt change in the external parameters. They correspond to Reeb chords connecting the initial and the terminal equilibrium Legendrian submanifolds.
One natural question that can be addressed by the tools of contact topology is the existence of an ultrafast processes that would connect the equilibria of two thermodynamic systems. As we discuss in Theorem 5.2 below, if two equilibrium thermodynamic Legendrians are related by a slow temperature non-decreasing global process, there is an ultrafast process which starts at an equilibrium state of the first system and ends at an equilibrium state of the second system.
Remark 1.1.
It should be mentioned that when the process is not quasi-static (that is, if we cannot assume that at each time moment the system is in the equilibrium), the temperature of the system is not well defined from the physical point of view. We always assume that the system exchanges heat with a thermostat (or reservoir) which has a well defined temperature, and for a system not in an equilibrium, we will consider the temperature of the reservoir at the given time. In an equilibrium, the temperature of the system is equal to the temperature of the thermostat (see [27], formula (44)). In physical terms, we tacitly assume that on the “boundary of the system” the temperature equals to that of the thermostat, while “inside” the temperature is defined only in an equilibrium. In this paper, we mostly consider the temperature as an external parameter and ignore its relaxation to the equilibrium. Elimination of the entropy and the temperature from the thermodynamic phase space is called reduction, see Sections 2.3 and 3.2.
We finish this brief introduction with Table 1 that sums up a contact topological glossary for thermodynamics.
Contact manifold (Reduced) thermodynamic phase space
Contact form The Gibbs fundamental form
Legendrian submanifold (Reduced) equilibrium submanifold
A non-negative path A non-equilibrium (temperature non-decreasing)
A non-negative path of Legendrians Slow (temperature non-decreasing) global process
Reeb chord between two Legendrians An instant relaxation in the thermodynamic limit
Table 1: Thermodynamic glossary
Organization of the paper. Section 2 discusses the construction of the equilibrium Legendrians starting with the microscopic models and passing to the thermodynamic phase space. We also introduce the reduction process that allows us to formally freeze certain intensive and extensive thermodynamic variables. In Section 3 we explain the relation between the first and second laws of thermodynamics and the positivity of the thermodynamic paths in phase space. The definitions of the slow, fast and ultrafast processes are discussed in Section 4. Section 5 uses some non-trivial results in symplectitc topology to obtain certain consequences of the existence of slow global processes connecting the equilibrium Legendrian submanifolds for two given thermodynamic systems. Finally, in Section 6, we discuss the ultrafast processes in the examples of the ideal gas and Stirling engine, and the Curie-Weiss magnet. Interestingly, for the Stirling engine, these ultrafast processes seem to be also observed experimentally [26].
Acknowledgment. We thank Shin-itiro Goto for very useful comments on the manuscript. Preliminary results of this paper were presented by L.P. at the Symplectic Geometry Seminar in IAS, Princeton. The recording of this talk, called “Contact topology meets thermodynamics”, is available on YouTube. ME was partially supported by the Israel Science Foundation grant 2033/23, LP was partially supported by the Israel Science Foundation grant 1102/20, and LR was partially supported by NSF grant DMS-2205497 and by ONR grant N00014-22-1-2174.
2 Thermodynamic equilibria as the Legendrian sub-manifolds
The contact geometric notions become relevant for thermodynamics when it is considered in the thermodynamic phase space – an odd dimensional space with the coordinates given by a thermodynamic potential and pairs of conjugate thermodynamic variables. It is equipped with the Gibbs fundamental form, which endows it with a contact structure. We shall deal with two versions of the thermodynamic phase space, the extended one described in Section 2.2, and the reduced one introduced in Section 2.3.
For reader’s convenience, we first recall in Section 2.1 the connection between the original microscopic description of thermodynamic equilibria and the macroscopic picture that eventually leads to the thermodynamic phase space. We follow the framework of [12], with some minor modifications.
2.1 Microscopic and macroscopic thermodynamic states
[…]
$S(\rho)=-\int_{M}\rho\ln\rho~{}d\mu,~{}~{}{\rho\in{\mathcal{P}}}.$ (1)
Note that $S(\rho)\geq 0$ for any $\rho\in{\mathcal{P}}$ by Jensen’s inequality.
The free energy of the system is
[…]
Here, $q\in{\mathbb{R}}^{n}$ is an exterior physical parameter and $T>0$ is the temperature of the system (see Remark 1.1). The Hamiltonian $H(q,m)$ often appears in the form
$H(q,m)=V_{int}(m)+V_{ext}(q,m),$ (3)
where $V_{int}(m)$ is the microscopic internal energy, and $V_{ext}(q,m)$ is related to an external influence on the system, as indicated by its dependence on the exterior parameter $q\in{\mathbb{R}}^{n}$. A special case of interest is when $V(q,m)$ is linear in $q$:
$V_{ext}(q,m)=(q\cdot\overline{V}(m)).$ (4)
Here, ()($\cdot$)( ⋅ ) denotes the standard inner product on ${\mathbb{R}}^{n}$, and $\overline{V}$ is a map from $M$ to ${\mathbb{R}}^{n}$.
For every $T>0$, and $q\in{\mathbb{R}}^{n}$ fixed, the functional $G(q,\cdot)$ is strictly convex on $\mathcal{P}$. A direct computation shows that its unique critical point, known as the Gibbs distribution, or an equilibrium distribution, is
[…]
Note that the entropy of a macroscopic state $\rho\in{\mathcal{P}}$ can be written as
$S(\rho)=-\frac{\partial{G}}{\partial{T}}(T,q,\rho).$ (6)
Similarly, for given a macroscopic state $\rho\in{\mathcal{P}}$ the generalized pressures [2] are defined as
[…]
Note that $p_{i}$ do not depend on $T$. The entropy and the generalized pressures are, in the sense of (6)-(7), the dual variables to the external physical parameters $T>0$ and $q\in{\mathbb{R}}^{n}$, respectively.
In the special case of the linear Hamiltonians as in (3)-(4), the pressures have the form
[…]
and free energy in (2) is
$G(T,q,\rho)$ $=U(\rho)-TS(\rho)-\sum_{j=1}^{n}p_{j}(\rho)q_{j}.~{}~{}q\in{\mathbb{R}}^{n},~{}\rho\in{\mathcal{P}},$ (8)
where $U(\rho)$, called the macroscopic internal energy, is defined by
[…]
2.2 A geometric interpretation of the equilibria and the extended thermodynamic phase space
These basic notions can be represented geometrically as follows. Given a macroscopic state $\rho\in{\mathcal{P}}$, we consider the point $(-G(T,q,\rho),S(\rho),T,p(q,\rho),q)$ as an element of ${\mathbb{R}}^{2n+3}$ with the coordinates $z=-G$, $S,T$, and $(p_{j},q_{j})$, $j=1,\ldots,n$. This space is called the extended thermodynamic phase space and is denoted by $\widehat{\mathcal{T}}$. We will refer to $T$ and $q=(q_{1},\ldots,q_{n})$ as the intensive variables, and to $S$ and $p=(p_{1},\ldots,p_{n})$ as the extensive variables.
The space $\widehat{\mathcal{T}}$ is naturally identified with the jet bundle
$J^{1}{\mathbb{R}}^{n+1}:={\mathbb{R}}(z)\times T^{*}{\mathbb{R}}^{n+1}\;,$
equipped with the canonical $1$-form
$\widehat{\lambda}=dz-SdT-\sum_{j=1}^{n}p_{j}dq_{j},$ (10)
called the Gibbs 1-form. It is a contact form: the maximal dimension of an integral submanifold $\Lambda$ of the distribution $\text{ker}(\lambda)$ equals $n+1$. Recall that $\Lambda$ is called integral if it is tangent to the distribution, that is, $\lambda$ vanishes on $T\Lambda$. The integral submanifolds of the maximal dimension $n+1$ are called Legendrian.
In our setting, a standard example of a Legendrian submanifold is the 1-jet of a smooth function $f:{\mathbb{R}}^{n+1}\to{\mathbb{R}}$,
$\Lambda_{f}=\Big{\{}(z,S,T,p,q,z)\in{{\mathbb{J}}^{1}{\mathbb{R}}^{n+1}}\Big{|}~{}z=f(T,q),~{}S=\frac{\partial{f}}{\partial{T}}(T,q),~{}p_{j}=\frac{\partial f}{\partial q_{j}}(T,q),j=1,\ldots,n\Big{\}}\;.$ (11)
More generally, given a smooth function
$\Psi:{\mathbb{R}}^{n+1}\times E\to{\mathbb{R}}\;,$
where $E$ is a space of auxiliary variables $\xi$, the set $\Lambda_{\Psi}$ of all $(z,S,T,p,q)\in{\mathbb{J}}^{1}{\mathbb{R}}^{n+1}$ such that there exists $\xi\in E$ so that
$z=\Psi(T,q,\xi),~{}\frac{\partial\Psi}{\partial\xi}(T,q,\xi)=0,~{}S=\frac{\partial\Psi}{\partial T}(T,q,\xi),~{}p=\frac{\partial\Psi}{\partial q}(T,q,\xi),$ (12)
is a (possibly singular) Legendrian submanifold for a generic $\Psi$. The function $\Psi$ is called a generating function of $\Lambda_{\Psi}$, and the variables $\xi$ are called the ghost variables.
A key observation is that (5)-(7) exactly mean that a Gibbs measure $\rho_{G}(T,q,\cdot)$ of a thermodynamic system corresponds to a point on the Legendrian submanifold $\Lambda$ of the form (12) with the elements $\rho\in{\mathcal{P}}$ being the ghost variables $\xi$. The generating function of $\Lambda$ is the (minus) free energy $-\Phi$ (we learned this observation from [23]). In other words, the equilibria of a thermodynamic system form a Legendrian submanifold $\widehat{\Lambda}$ in $\widehat{\mathcal{T}}$, as we vary the external parameters $T$ and $q$, that themselves are coordinates on $\widehat{\mathcal{T}}$.
2.3 The reduced thermodynamic phase space
The aforementioned Legendrian submanifold $\widehat{\Lambda}$ includes all possible equilibria, for all values of $T>0$ and $q\in{\mathbb{R}}^{n}$. As the physical parameters change, one can, in principle, think of time-dependent thermodynamic non-equilibrium processes that start and end at two corresponding equilibria as paths connecting two points on $\widehat{\Lambda}$. This, however, is somewhat unwieldy, and it is convenient to work in the reduced thermodynamic phase space. With an eye toward applications, we present a rather general reduction procedure. Fix $1<k\leq\ell\leq n$ and set
$I=\{k+1,\dots,\ell\},~{}~{}E=\{\ell+1,\dots,n\}.$
Consider the affine subspace $h\subset J^{1}{\mathbb{R}}^{n+1}$ obtained by fixing the values of the temperature and the intensive variables with indices from $I$, and setting the extensive variables with indices from $E$ to be zero:
$h=\{T=T^{0};q_{i}=q_{i}^{0},\;i\in I;\;p_{e}=0,\;e\in E\}\;.$ (13)
This is an $n+2+k$ dimensional sub-space. The reduced thermodynamic phase space $\mathcal{T}$ is obtained by the projection of $h$ along the directions corresponding to the variables $S$, $p_{i},i\in I$ and $q_{e},e\in E$, and is $12k+$ 1 dimensional.
The reduced thermodynamic phase space is naturally identified with $J^{1}{\mathbb{R}}^{k}$. Consider now the projection $J^{1}{\mathbb{R}}^{n+1}\to J^{1}{\mathbb{R}}^{k}$,
$(z,S,T,p_{1},q_{1},\dots,p_{n},q_{n})\mapsto(z,p_{1},q_{1},\dots,p_{k},q_{k})\;,$
and let
$\lambda=dz-\sum_{j=1}^{k}p_{j}dq_{j},$ (14)
be the standard contact structure on $J^{1}{\mathbb{R}}^{k}$. Given an equilibrium Legendrian submanifold $\widehat{\Lambda}\subset\widehat{\mathcal{T}}$, its reduction $\Lambda$ is defined as the image of $\widehat{\Lambda}\cap h$ under the above-mentioned projection. It follows from (10) and the definition (13) of $h$ that, under certain transversality assumptions, the set $\Lambda$ is also a Legendrian submanifold of $J^{1}{\mathbb{R}}^{k}$ with respect to the standard contact structure. Physically, $\Lambda$ represents the collection of all equilibria of the system with the temperature $T$ and the intensive variables $q_{i}$, $i\in I$ fixed and $p_{e}=0$, $e\in E$.
Enlarging the original system by adding unspecified thermodynamic variables (for example, considering a system of pair-wise interacting Curie-Weiss magnets discussed in Example 3.3 below) and reducing them at some fixed values will lead to more complicated families of Legendrians which could be considered as perturbations of $\Lambda$. In fact, one can get in this way every Legendrian which is obtained from the zero section by a compactly supported Hamiltonian isotopy. This fact and its extension to jet bundles of manifolds follows from the theory of generating functions, (see e.g. [7] for a detailed discussion). It would be interesting to elaborate meaningful examples of such perturbations.
Passing to the next level of abstraction, we can assume that the space of intensive variables is a smooth manifold $X$ and the thermodynamic phase space (either extended, or reduced) is $J^{1}X={\mathbb{R}}(z)\times T^{*}X(p,q)$ equipped with the contact form $\lambda=dz-pdq$. The submanifolds $\Lambda$ consisting of all equilibria of such system are Legendrian: $\dim\Lambda=\dim X$ and $\lambda|_{T\Lambda}=0$.
Remark 2.1.
In the context of thermodynamics, there is a discrepancy between mathematical theory which requires a reasonable behavior “at infinity” of the Legendrian submanifolds and the pool of physical examples. This resembles the situation in classical mechanics where the best theory exists for the compactly supported case, while physically meaningful Hamiltonians are not compactly supported. Nevertheless, sometimes one can close this gap by topological tricks, see e.g. the proof of Theorem 6.4 below for a reduction of a physical problem to $J^{1}S^{1}$.
3 Positivity of the thermodynamic paths
In this section, we will see that the laws of thermodynamics imply that the paths that correspond to physical processes must be non-negative with respect to the corresponding contact forms both in the extended and reduced thermodynamic phase spaces, a minor modification of a result of [20].
3.1 The thermodynamic paths in the extended thermodynamic phase space
We will now consider, for the sake of convenience, thermodynamic systems with affine Hamiltonians of the form (3)-(4) so that the free energy has the form (8). So far, we have described the collections of thermodynamic equilibria as points on the Legendrian submanifolds, either in the full thermodynamic phase space or in its reduced version. A thermodynamic path is a smooth path
$\gamma(t)=(-G(T(t),q(t),\rho(t)),S(\rho(t)),T(t),p(q(t),\rho(t)),q(t))\subset\widehat{\mathcal{T}},$ (15)
in the phase space generated by a smooth time-dependent process, with $T(t)>0$, $q(t)\in{\mathbb{R}}^{n}$ and $\rho(t)\in{{\mathcal{P}}}$, that may be out-of-the-equilibrium. As we have already mentioned in Remark 1.1, we always assume that the system exchanges heat with a thermostat (or reservoir) which has a well defined temperature, and the coordinate $T(t)$ of the path $\gamma$ is the temperature of the reservoir at the time $t$.
A reformulation of an important observation of [20] is that the laws of thermodynamics impose a constraint on the paths in the thermodynamic phase that can be physically realized. To this end, let us recall the following
Definition 3.1.
A path $\gamma(t)\subset J^{1}{\mathbb{R}}^{n+1}$ is called non-negative if $\widehat{\lambda}(\dot{\gamma}(t))\geq 0$ for every $t$.
To derive the aforementioned constraint, let $(T(t),q(t)),\rho(t))$ be a smooth path and $\dot{\gamma}(t)$ and $\dot{\rho}(t)$ be, respectively, the tangent vectors to the corresponding thermodynamic path $\gamma(t)$ in (15) and to $\rho(t)\in{\mathcal{P}}$. By the first law of thermodynamics, the infinitesimal increment of the macroscopic internal energy defined in (9) can be written as the difference
$dU(\dot{\rho})=\Delta Q-\Delta W(\dot{\gamma})\;.$ (16)
Here, $\Delta Q$ is the infinitesimal amount of heat supplied to the system, and
$\Delta W(\dot{\gamma})=-\sum_{j=1}^{n}q_{j}dp_{j}(\dot{\gamma})$ (17)
is the work done by the system (see [22]). In fact Jarzynski in [22] distinguishes between two types of work, inclusive and exclusive, and the one given in (17) is the exclusive one. In addition, by the second law of thermodynamics, the increment of the entropy is
$dS(\dot{\gamma})=\frac{\Delta Q}{T}+d_{irr}S(\dot{\gamma}).$ (18)
Here, $d_{irr}S(\dot{\gamma})$ is the irreversible entropy production, see [8], equation (3.8). The second law of thermodynamics states that $d_{irr}S(\dot{\gamma})\geq 0$ for all $t>0$. If $d_{irr}S(\dot{\gamma})=0$ for all $t$, the process is reversible. Otherwise, it is called irreversible. Combining (16), (17) and (18), together with the expression (8) for the free energy and the definition (10) of $\widehat{\lambda}$, we get
$Td_{irr}S(\dot{\gamma})=-dU(\dot{\rho})+TdS(\dot{\gamma})+\sum_{j=1}^{n}q_{j}dp_{j}(\dot{\gamma})=\widehat{\lambda}(\dot{\gamma})\;.$
We conclude from the second law of thermodynamics that
[…]
an observation that can be found in a slightly different form in [20]. Note that for a reversible process we always have $\widehat{\lambda}(\dot{\gamma})=0$. This is closely related to the fact that the set of equilibria of a system forms a Legendrian submanifold.
3.2 The thermodynamic paths in the reduced thermodynamic phase space
We now discuss the non-negativity of the thermodynamic paths in the reduced thermodynamic phase space. Suppose that the reduction is made with respect to the intensive variables with the indices from $I$ and the extensive variables with the indices from $E$, where
$I=\{k+1,\dots,\ell\},~{}~{}E=\{\ell+1,\dots,n\}.$
Consider the projection $J^{1}{\mathbb{R}}^{n+1}\to J^{1}{\mathbb{R}}^{k}$,
$(z,S,T,p_{1},q_{1},\dots,p_{n},q_{n})\mapsto(z,p_{1},q_{1},\dots,p_{k},q_{k})\;,$
and let $\lambda$ be the standard contact structure on $J^{1}{\mathbb{R}}^{k}$, as defined in (14). Take any thermodynamic path
$\widehat{\gamma}(t)=(z(t),S(t),T(t),p(t),q(t))\in\widehat{\cal T}.$
Assumption 3.2.
While considering thermodynamic processes given by a path in the reduced thermodynamic phase space, we assume that
$A:=S(t)\dot{T}(t)+\sum_{j\in I}p_{j}(t)\dot{q}_{j}(t)\geq 0\;.$ (20)
Furthermore, we assume that $p_{e}=0$, $e\in E$.
Let us illustrate inequality (20) in some situations. Generally, if the temperature $T$ is the only reduced variable, then $A\geq 0$ whenever the temperature is non-decreasing. This is because we always have $S(t)\geq 0$ by the definition (1) of the entropy. Similarly, in the ideal gas example considered in Section 6.1 below, the volume $V$ and the negative pressure $-P$ are pair-wise conjugate extensive and intensive variables, respectively, and the volume is, of course, always non-negative. Thus, if we reduce the pressure only, the assumption above reads that the pressure is non-increasing. On the other hand, in the Curie-Weiss model discussed in Example 3.3 and Section 6.2 below, the magnetization $M$ (dual to the external magnetic field $H$ that is an intensive variable) is an extensive variable that is not necessarily positive. In that context, when the external magnetic field is reduced, assumption (20) does not directly translate into its monotonicity in time.
A more general interpretation of (20) comes from the fact that in an equilibrium it can be written as
$A=-\frac{\partial G}{\partial T}\dot{T}-\sum_{j\in I}\frac{\partial G}{\partial q_{j}}\dot{q_{j}}\;.$ (21)
Thus, the first term in the right side of (20) is the rate of change of the free energy due to the change in the temperature, while the second term is, in the terminology of [22], the inclusive work done by the system. In other words $A$ is the decrement of the free energy due to the change of the reduced intensive variables from $I$. We shall tacitly assume that the same interpretation holds for non-equilibrium processes which are close to equilibrium ones (e.g., for quasi-static processes considered in Section 4 below). For such processes the first part of Assumption 3.2 can be restated as follows:
Throughout the process, the reduced variables contribute either a decrease (22)
or no change to the free energy.
Let
$\gamma(t)=(z(t),p_{1}(t),q_{1}(t),\dots,p_{k}(t),q_{k}(t)),$
be the projection of $\widehat{\gamma}(t)$ on the reduced thermodynamic phase space. Note that by (19), the path $\widehat{\gamma}$ is non-negative in the extended thermodynamic phase space $\widehat{\cal T}$. As $p_{e}=0$ for $e\in E$, we have
$0\leq\hat{\lambda}\Big{(}\frac{d\widehat{\gamma}}{dt}(t)\Big{)}$ $=\dot{z}(t)-S(t)\dot{T}(t)-\sum_{i=j}^{n}p_{j}(t)\dot{q}_{j}(t)$ (23)
$=\dot{z}(t)-S(t)\dot{T}(t)-\sum_{j=1}^{k}p_{j}(t)\dot{q}_{j}(t)-\sum_{j=k+1}^{l}p_{j}(t)\dot{q}_{j}(t).$
We deduce from the aforementioned assumptions on the path $\widehat{\gamma}(t)$ that
$\lambda(\dot{\gamma}(t))=\dot{z}(t)-\sum_{j=1}^{k}p_{j}(t)\dot{q}_{j}(t)\geq S(t)\dot{T}(t)+\sum_{j=k+1}^{l}p_{j}(t)\dot{q}_{j}(t)\geq 0.$ (24)
Thus, the projection $\gamma(t)$ of $\widehat{\gamma}(t)$ is again non-negative with respect to the standard contact structure $\lambda$. We call $\gamma$ a reduced thermodynamic path.
In the abstract setting of a thermodynamic system where the space of the intensive variables is a manifold, as discussed above Remark 2.1, admissible thermodynamic paths are given by non-negative paths $\gamma(t)$, i.e.,
$\lambda(\dot{\gamma}(t))\geq 0\;.$ (25)
Example 3.3.
(The Curie-Weiss magnet) Let $n=k=1$, $p_{1}=M$ and $q_{1}=H$ be the magnetization and the external magnetic field of a mean-field Ising model (a.k.a. the Curie-Weiss magnet) in the thermodynamic limit. Here, the temperature $T$ and the external magnetic field $H$ are the intensive variables, and the entropy $S$ and the magnetization $M$ are the extensive variables. We assume that the system is subject to the background magnetic field $H_{back}$ playing the role of a parameter (see Remark 6.1 below). The total magnetic field acting on the magnet is then $H+H_{back}$. The equilibrium Legendrian $\widehat{\Lambda}$ in $J^{1}{\mathbb{R}}^{2}$ is given by the equations (see Section 2 in [17])
$z=T\ln\left(2\cosh\frac{H+H_{back}+bM}{T}\right)-\frac{b}{2}M^{2},\;M=\tanh\left(\frac{H+H_{back}+bM}{T}\right)\;,$ (26)
$S=-\frac{1-M}{2}\ln\frac{1-M}{2}-\frac{1+M}{2}\ln\frac{1+M}{2}\;.$
Here, $b>0$ is the spin interaction parameter. The reduced Legendrian $\Lambda\subset J^{1}{\mathbb{R}}$ (where we only reduce the temperature and the entropy) is given by the first two equations and depends on $T$ and $H_{back}$ as parameters. We sometimes emphasize this by writing $\Lambda=\Lambda(T,H_{back})$. Given a thermodynamic temperature-non-decreasing process in $\widehat{\mathcal{T}}=J^{1}{\mathbb{R}}^{2},$ satisfying (19), its reduction is a non-negative path in $\mathcal{T}=J^{1}{\mathbb{R}}$.
4 Classification of thermodynamic processes
As we have already stated in the introduction, the main theme of this note is modeling of the non-equilibrium thermodynamic processes. We distinguish between three types of irreversible processes in the reduced (at least with respect to temperature) thermodynamic phase space. We adopt Assumption 3.2.
I. Slow (quasi-static) global processes: They are given by non-negative families (or paths) of Legendrian (equilibrium) submanifolds. Mathematically, this means that we have an initial Legendrian submanifold $\Lambda\subset\mathcal{T}$ and a map
$\Gamma:\Lambda\times[0,\tau]\to\mathcal{T},\;\Gamma(x,0)=x\;\;\forall x\in\Lambda,$
such that for each $t$, the image of $\Gamma(\cdot,t)$ is a Legendrian submanifold – such a map $\Gamma$ is called a (parameterized) Legendrian isotopy, see Section 5 below222 As we already stated in the introduction, by a global process we mean a process defined at all macroscopic equilibrium states of the system. In the geometric language, this means that the map $\Gamma$ describing this process is defined on the direct product of the whole equilibrium Legendrian submanifold $\Lambda$ – and not only of its part – with the time interval. All Legendrian submanifolds that we consider are assumed to be closed subsets of the thermodynamic phase space.. Furthermore, we assume that each individual path $t\mapsto\Gamma(x,t)$ (for a fixed $x$) is non-negative, that is, a thermodynamic path satisfying (19). (Such a Legendrian isotopy is called non-negative, see Section 5 below). Physically, this means that the change is so slow that, at each moment in time, the system is in equilibrium. An example is provided by the Curie-Weiss magnet whose temperature is slowly increasing, yielding a path of submanifolds $\Lambda(T,0)$ from Example 3.3.
A thermodynamic path $t\mapsto\Gamma(x,t)$ appearing in a slow (quasi-static) global process $\Gamma$ will be called a slow thermodynamic path.
II. Fast processes: Here we consider a fast [^3] Here and below ’fast’ means faster than quasi-static and slower than instantaneous. The convergence to the equilibrium is exponential as governed by the first positive eigenvalue of the Fokker-Planck operator. evolution of an individual macroscopic state. (In physical terms, one can think of this evolution as a relaxation of a thermodynamic system in a particular macroscopic state after an abrupt change in the extensive variables – see e.g. [12] for a discussion of such a relaxation of a Curie-Weiss magnet after a sudden change of its temperature and the exterior magnetic field). The evolution is given by a thermodynamic path $\gamma(t)=(z(t),p(t),q)$, in the reduced thermodynamic phase space, generated by an evolution of the temperature $T(t)\geq 0$ and of the probability density $\rho(t)\in{\mathcal{P}}$. We consider the temperature $T(t)$ as a prescribed non-decreasing function, with given $T_{0}=T(0)$ and $T_{\infty}=\lim_{t\to+\infty}T(t)$. The intensive variables $q$ remain fixed. The evolution of $\rho(t)$ is described by the Fokker-Plank equation
$\dot{\rho}=-\nabla_{\rho}G,$ (27)
that involves $T(t)$ as a coefficient, see [12], where such models are elaborated for time-independent temperature. The corresponding path $\gamma(t)=(z(t),p(t),q)$, with $z=-G(T,q,\rho)$, converges to the equilibrium submanifold corresponding to the terminal temperature $T_{\infty}$ of the process.
We claim that $\gamma(t)$ necessarily non-negative. Indeed, recalling (6) and since $q$ is kept fixed, we get
$\lambda(\dot{\gamma})=\dot{z}=d_{\rho}G(\nabla_{\rho}G)+S\dot{T}\geq 0\;,$ (28)
and the claim follows.
An example is given by the thermodynamic path corresponding to the transition from the metastable to the stable equilibrium of the Curie-Weiss magnet, see [12] for a discussion of metastability and further references.
III. Ultrafast (instantaneous) processes: They correspond to Reeb chords, i.e., to pieces of trajectories of the Reeb vector field $\partial/\partial z$ connecting the initial and the terminal equilibrium Legendrian submanifolds; see [12] for a detailed discussion as well as Section 6 below. Both variables $(p,q)$ remain constant along such a process, while $z$ is increasing.
Remark 4.1.
As explained in [12], if the endpoint of the Reeb chord corresponds to an unstable or metastable equilibrium of the terminal system, there will be further relaxation to the stable equilibrium. We do not address the issue of the stability of equilibria in the context of Reeb chords, and it would be interesting to understand if contact topology can be helpful here. For simplicity, we stick to the terminology in III.
Remark 4.2.
The Reeb chords represent instant jumps from its initial point to the terminal point. In contrast to the slow and fast processes, the intermediate points of the chord seem to have no physical meaning.
Remark 4.3.
While ultrafast processes correspond to Reeb chords, the reverse is not true: Reeb chords can correspond to fast processes and may also appear as paths of individual states in slow global processes. This is illustrated in Example 6.3 below.
5 Slow global processes and partial order on Legendrians
Looking at an equilibrium Legendrian submanifold $L$ of $\mathcal{T}$, the reduced thermodynamic space with the temperature and the entropy being the only reduced variables, we cannot, in general, reconstruct the temperature unless we have an additional knowledge about the system. However, the existence of a temperature-nondecreasing slow global process which connects two given properly embedded Legendrians $L_{0}$ and $L_{1}$ imposes a contact topological constraint on $L_{0}$ and $L_{1}$ which we explain in this section.
All Legendrians considered below are assumed to be properly embedded, i.e. their intersection with every compact subset of $\mathcal{T}$ is compact.
Let us pass for a moment to a more abstract setting of a contact manifold $\Sigma$ with a co-oriented contact structure $\xi=\text{ker}(\alpha)$. Here, $\alpha$ is a contact form on $\Sigma$ defining the co-orientation. Let $\Lambda$ be a properly embedded Legendrian submanifold. A parameterized Legendrian isotopy of $\Lambda$ is a map
$\Gamma:\Lambda\times[0,1]\to\Sigma,\;\Gamma(x,0)=x\;\;\forall x\in\Lambda,$
such that each $\Lambda_{t}:=\Gamma(\Lambda\times t)$ is a Legendrian submanifold. Let
$v_{t}(x):=\frac{\partial\Gamma}{\partial t}(x,t),~{}~{}(x,t)\in\Lambda\times[0,1],$
be the vector field of the isotopy. We say that $\Gamma$ is compactly supported if so is the (time-dependent) vector field $v_{t}$. We call $\{\Lambda_{t}\}$, $t\in[0,1]$, a non-parameterized Legendrian isotopy, or just a Legendrian isotopy, and $\Gamma$ its parameterization. We say that a Legendrian isotopy is compactly supported if it admits a compactly supported parameterization.
Let $[\Lambda]$ be the family of all Legendrians in $(\Sigma,\xi)$ that can be obtained from a given properly embedded Legendrian $\Lambda$ by a compactly supported Legendrian isotopy. We say that a compactly supported Legendrian isotopy $\Lambda_{t}$, $t\in[0,1]$, of Legendrians in $[\Lambda]$ is non-negative if it admits a parameterization with a vector field $v_{t}$ satisfying $\alpha(v_{t})\geq 0$ (note that if this holds for one parameterization, then it holds for all the others). We write $\Lambda_{0}\preceq\Lambda_{1}$ if there exists a non-negative Legendrian isotopy from $\Lambda_{0}$ to $\Lambda_{1}$.
With the geometric terminology as above, slow global thermodynamic processes are described by compactly supported parameterized Legendrian isotopies in the thermodynamic phase space $\Sigma:=J^{1}X$ equipped with the contact structure as above. We refer to the slow global processes satisfying Assumption 3.2 as the admissible global processes. Admissible processes correspond to non-negative isotopies. Recall that when the temperature and the entropy are the only reduced variables, admissibility means that the temperature is non-decreasing.
We are now ready to present a few thermodynamic consequences of contact topology. The topological results used below are established in the literature for closed manifolds and are, in part, folklore for open ones. After stating them, together with their thermodynamic interpretations, we outline their proofs and provide relevant references.
The class $[\Lambda]$ is called orderable if the binary relation $\preceq$ is a partial order. A Reeb chord between two Legendrians is called non-trivial if it has positive length (as opposed to a point). The space of intensive variables $X$ below is assumed to be a connected manifold without boundary (either closed, i.e. compact, or open, i.e., non-compact).
Theorem 5.1.
For any (open or closed) smooth connected manifold $X$, the class $[0_{X}]$ of the zero section $0_{X}$ in $J^{1}X=T^{*}X\times{\mathbb{R}}$ is orderable. If two distinct equilibrium Legendrians $\Lambda_{0},\Lambda_{1}\in[0_{X}]$ are related by the partial order, $\Lambda_{0}\preceq\Lambda_{1}$, there is no admissible global process taking $\Lambda_{1}$ to $\Lambda_{0}$.
The next two results establish a connection between global slow and ultrafast processes.
Theorem 5.2.
If two distinct Legendrians $\Lambda_{0},\Lambda_{1}\in[0_{X}]$ satisfy $\Lambda_{0}\preceq\Lambda_{1}$, there exists a non-trivial Reeb chord starting on $\Lambda_{0}$ and ending on $\Lambda_{1}$. In particular, if two equilibrium thermodynamic Legendrians from $[0_{X}]$ are related by an admissible global process, then there is an ultrafast process which starts at an equilibrium state of the first system and ends at an equilibrium state of the second system.
Theorem 5.3.
For every two distinct Legendrians in $[0_{X}]$ there exists a non-trivial Reeb chord connecting them (in an unspecified order). If these are the equilibrium Legendrians of two thermodynamic systems, there exists an ultrafast process that starts at an equilibrium state of one system and ends at an equilibrium state of the other.
Comments on the proofs:
Case I: $X$ is a closed manifold.
In this case, Theorem 5.1 follows from [5, Proposition 5], [7, Theorem 1]. Theorem 5.2 can be deduced from the theory of generating functions and their spectral selectors [6, 7, 29] along the following lines. Assume without loss of generality that $\Lambda_{0}$ is the zero section. Let $c_{+}$ and $c_{-}$ be the spectral selectors corresponding to the fundamental class (i.e., the class of $X$) and the class of the point in the homology $H_{*}(X,\mathbb{Z}_{2})$ with the $Z_{2}$-coefficients, respectively. If $\Lambda_{0}\preceq\Lambda_{1}$, then $c_{+}(\Lambda_{1})\geq 0$. Since $\Lambda_{0}\neq\Lambda_{1}$, we have $c_{+}(\Lambda_{1})>0$. It follows from the spectrality property of the spectral selectors (see e.g. Proposition 2.4(i) in [29]) that there exists a chord of time-length $c_{+}(\lambda_{1})$ starting on $\Lambda_{0}$ and ending on $\Lambda_{1}$, as required. Theorem 5.3 follows from the fact that since $\Lambda_{0}\neq\Lambda_{1}$, at least one of $c_{-}(\Lambda_{1})$ or $c_{+}(\Lambda_{1})$ is distinct from 00.
A different argument which deduces Theorems 5.2 and 5.3 from the orderability of the class $[0_{X}]$ is discovered in a recent paper by Allais and Arlove [1], see Corollary 1.5 and Theorem 4.7 therein. Interestingly enough, this argument works on arbitrary contact manifolds, with $[0_{X}]$ being replaced by arbitrary orderable Legendrian isotopy class of a closed Legendrian submanifold. Our exposition is influenced by this paper.
Case II: $X$ is an open manifold.
When $X={\mathbb{R}}^{n}$, all the results follow from the spectral selectors constructed in [3]. Existence of generating functions and the first properties of spectral selectors on general open manifold is elaborated in [6]. A quick reduction of the open case to the closed one is obtained by the following construction. Choose an exhausting proper bounded from below function $f:X\to{\mathbb{R}}$. Set $X_{C}:=\{f\leq C\}$, where $C$ is a regular value of $f$. Let $X_{C}^{\circ}$ be the interior of $X_{C}$. Any compactly supported Legendrian isotopy in $J^{1}X$ is supported in $J^{1}X_{C}^{\circ}$ for some large $C$. Now do the following trick: pass to the double of $X_{C}$, denoted by $Y_{C}$ – in other words, glue $X_{C}$ with itself along the boundary $\{f=C\}$. Crucially, $Y_{C}$ is a closed manifold.
Theorem 5.1 follows from the fact that every non-negative Legendrian loop in $J^{1}X$ corresponds to such a loop in $J^{1}Y_{C}$ and hence is constant by Theorem 5.1 applied to $Y_{C}$. Theorems 5.2 and 5.3 follow from the fact that in the closed case the chords constructed in Step 1 are non-trivial. Thus, when we apply these theorems to $J^{1}Y_{C}$, we get the chords lying in $J^{1}X_{C}$. This completes the proof.
Remark 5.4.
We have considered two models of thermodynamic processes: thermodynamic paths and paths of Legendrian submanifolds, both in the thermodynamic phase space. It is natural to consider a model which interpolates between these two, as follows. Given an initial Legendrian submanifold $\Lambda_{0}\subset\mathcal{T}$, we consider a smooth (not necessarily Legendrian!) isotopy
$\Gamma:\Lambda_{0}\times[0,1]\to\mathcal{T},\;\Gamma(x,0)=x\;\;\forall x\in\Lambda,$
such that $\Lambda_{1}:=\Gamma(\Lambda_{0}\times 1)$ is a also Legendrian submanifold (but intermediate $\Lambda_{t}:=\Gamma(\Lambda\times t)$ do not have to be Legendrian). Additionally, we assume that all individual paths $t\mapsto\Gamma(x,t)$ are non-negative (or, as a variation of this problem, positive, i.e., transversal to the contact hyperplanes and respecting the coorientation). We call such smooth isotopies non-negative/positive. In other words, we model a process by a path of submanifolds connecting the initial and the terminal equilibrium submanifolds without assuming that in the course of the process the system is always in an equilibrium, but insisting that the individual trajectory of every state agrees with the laws of thermodynamics, as in (19) and (24). At the moment, it is not clear to which extent the non-negativity/positivity assumption is restrictive in this context. On one hand, there do exist examples of pairs of (compact) Legendrian submanifolds that can be connected by a non-negative smooth isotopy but cannot be connected by a non-negative Legendrian isotopy. For instance, it is not hard to construct a non-negative smooth isotopy between $\{z=1,p=0\}$ and $\{z=-1,p=0\}$ in $J^{1}S^{1}$, though there is no non-negative Legendrian isotopy between these curves. On the other hand, it is unclear how to characterize or detect such pairs in general.
Remark 5.5.
We refer to [24] for appearance of some other order relations in thermodynamics. It would be interesting to relate the discussion in [24] with the partial order on Legendrians considered above.
6 Ultrafast processes as Reeb chords
In this section, we discuss two examples of ultrafast processes: an ideal gas and the Curie-Weiss magnet. The ideal gas model is directly related to the Stirling engine, a well known device, where the ultrafast process can be observed experimentally [26]. In the Curie-Weiss magnet context, we also establish the existence of an ultrafast process connecting an equilibrium of a magnet at a given temperature to a perturbation of the magnet at a larger temperature.
6.1 An isochoric temperature and pressure jump
Consider the lattice gas consisting of $N$ particles occupying $V$ sites at temperature $T$ in the thermodynamic limit $V,N\to+\infty$ so that $V/N\to v$. Furthermore, we assume that the gas is diluted, $v\gg 1$. This is a model approximating the ideal gas – see [13, Section 1.3.2] for details.
In this model, the temperature $T$ and the pressure $P$ are the intensive variables, while the volume $v$ is the extensive variable. We will not need to use the entropy in this example. If the Bolzmann and the gas constants are taken to be $1$, in an equilibrium we have
$(P+P_{back})v=T.$ (29)
Here, $P_{back}$ is a background pressure that we treat as a physical parameter of the system. If the molecules do not interact so that the internal energy vanishes, then the Hamiltonian in this model is linear, of the form (3)-(4) with $V_{int}(m)=0$. Then, a straightforward calculation following [13] shows that the equilibrium free energy $G$ in the ideal gas approximation is
$G=-T\ln v=T\ln((P+P_{back})/T).$
We warn the reader that in this model the internal energy per particle vanishes, as opposed to the standard model of the ideal gas where it equals $\frac{3}{2}T/v$.
Assume now that the temperature jumped from $T_{0}$ to $T_{1}>T_{0}$ and the background pressure simultaneously jumped from $P_{back}=0$ to $P_{back}=c>0$, while the volume is kept fixed. One can think that the gas is contained in a chamber with one of the walls being a piston of area $a$, and one instantly changes the temperature and simultaneously switches on a force $f$ applied to the piston and directed towards the gas so that $c=f/a$.
Given the parameters $T_{0}$, $T_{1}$ and $c$, there exist unique values of the volume $v$ and the initial pressure $P_{0}$ when such a jump corresponds to the ultrafast process, see Figure 1. This happens if, after the jump, the system lands in the equilibrium of the terminal system, so that
$v=\frac{T_{0}}{P_{0}}=\frac{T_{1}}{P_{0}+c}\;.$ (30)
We compute that
$P_{0}=\frac{cT_{0}}{T_{1}-T_{0}},\;v=\frac{T_{1}-T_{0}}{c}\;.$ (31)
In particular, if the background pressure jump $c\ll 1$ is small, and $T_{1}-T_{0}=O(1)$, the volume $v$ is large in accordance to the ideal gas approximation.
Figure 1: $T_{0}=1,T_{1}=5,c=2$
Note that the change in the Gibbs free energy is negative if the pressure jump $c<T_{1}-T_{0}$:
$G_{1}-G_{0}=(T_{0}-T_{1})\ln v<0\;.$ (32)
Let us mention that isochoric temperature and pressure jumps appear as two of the four segments of the Stirling engine cycle [^4] We thank S. Goto for the reference to the Stirling engine. For a concise description of the Stirling engine see https://en.wikipedia.org/wiki/Stirling_engine. The attribution for Figure 2 is to Cristian Quinzacara, CC BY-SA 4.0 https://creativecommons.org/licenses/by-sa/4.0, via Wikimedia Commons. as depicted in Figure 2. It is tempting to view the segment of the engine where the temperature goes up at a fixed volume as precisely the above Reed chord, since in its derivation we have only used the ideal gas equilibrium law $PV=T$ that also holds for the Stirling engine. We shall see below that this is indeed the case, albeit after the shift of the terminal Legendrian submanifold along the $P$-axis by a suitable background pressure. Let us also mention that a recent paper [26] dealing with a microscale Stirling engine states that “the isochoric transitions were nearly instantaneous and occurred on millisecond time scales”.
Figure 2: The Stirling engine cycle, with $T_{0}=T_{C}<T_{H}=T_{1}$.
Remark 6.1.
When discussing an instantaneous isochoric process, as in the Stirling engine, temperature acts as a parameter, while pressure is a macroscopic variable. Introducing an artificial parameter, the background pressure, such that the total pressure is given by $P+P_{back}$, allows us to describe the simultaneous jump of two parameters: $T$ jumps from $T_{0}$ to $T_{1}$, and $P_{back}$ from 00 to $c$. More generally, this trick applies to processes in which intensive variables exhibit instantaneous jumps while the corresponding extensive variables remain unchanged, see Section 6.2 for another example. This not only streamlines the exposition, but, more importantly, gives rise to the connection between such processes and Reeb chords.
Let us explain now how such ultrafast process corresponds to a Reeb chord between two Legendrian submanifolds. We work in the reduced thermodynamic phase space $(z,p,q)$ where $z=-G,p=v,q=-P$. This notation is in accordance with the fact that volume $v$ is an extensive variable and the pressure $P$ is the intensive variable [^5] We hope that the difference between $p$ and $P$ will not cause confusion for the reader..
The Legendrian submanifold corresponding to the ideal gas is given by
$\Lambda(T,P_{back})=\{(z,p,q)\in{\mathbb{R}}^{3}\;:\;z=\phi_{T}(q-P_{back}),\;p=\phi^{\prime}_{T}(q-P_{back})\}\;,$ (33)
where $\phi_{T}(q)=-T\ln(-q/T)$. Let us set $\Lambda_{0}=\Lambda(T_{0},0)$ and $\Lambda_{1}=\Lambda(T_{1},c)$.
To detect the chord connecting $\Lambda_{0}$ and $\Lambda_{1}$, one needs to solve the equation
$p_{0}=p_{1}=-T_{0}/q=-T_{1}/(q-c),$
which coincides with (30). Thus, the ultrafast process discussed above, corresponds exactly to the Reeb chord connecting $\Lambda_{0}$ with $\Lambda_{1}$. Furthermore, the orientation of the chord is given by $\partial/\partial z$ since $z_{1}-z_{0}>0$ by (32) and, as we recall, $z=-G$.
Let us illustrate the Reeb chord after the change of variables
$\overline{z}=z-\phi_{T_{0}}(q),~{}~{}\overline{p}=p-\phi^{\prime}_{T_{0}}(q),~{}~{}\overline{q}=q\;.$
It preserves the Gibbs form, sends $\Lambda_{0}$ to the zero section $\overline{\Lambda}_{0}$, and $\Lambda_{1}$ to
$\overline{\Lambda}_{1}=\{(z,p,q)\in{\mathbb{R}}^{3}\;:~{}\overline{z}=\psi(\overline{q}),~{}\overline{p}=\psi^{\prime}(\overline{q})\},\;$
where
$\psi(\overline{q})=\phi_{T_{1}}(\overline{q}-c)-\phi_{T_{0}}(\overline{q})\;.$
The projections of the resulting Legendrians and the chord onto the $(q,z)$ plane (the front projection) are given on Fig. 3.
Figure 3: $T_{0}=1,T_{1}=5,c=2$
Remark 6.2.
It sounds likely that there is no non-negative Legendrian isotopy (that is, a slow global thermodynamic process) connecting $\Lambda_{0}$ and $\Lambda_{1}$. To make this statement rigorous, one has to find a way to deal with the lack of compactness in this example, cf. Remark 2.1 above. For instance, let us modify $\overline{\Lambda}_{1}$ below $\overline{\Lambda}_{0}$ so that it coincides with the line $\{z=-1,p=0\}$ outside a compact set. Twist now $J^{1}{\mathbb{R}}$ into $J^{1}S^{1}$ by taking the quotient by $(z,p,q)\mapsto(z,p,q+C)$ with a large constant $C$, so that $\overline{\Lambda}_{0}$ and $\overline{\Lambda}_{1}$ are modified to the Legendrian circles $\Lambda^{\prime}_{0}$ and $\Lambda^{\prime}_{1}$. One can show that these circles are incomparable with respect to the partial order. This can be proved with the help of Legendrian spectral invariants which are monotone with respect to non-negative Legendrian isotopies, see [7, Section 4.1].
Example 6.3.
Here, we illustrate that Reeb chords can also correspond to a slow thermodynamic path, still on the ideal gas example. Suppose the temperature increases slowly with time $0\leq t\leq 1$, given by $T(t)=T_{0}+(T_{1}-T_{0})t$, and the external pressure also increases slowly $P_{ext}(t)=ct$, while the volume remains constant, $v(t)=v$. The Lagrangian projection of the equilibrium Legendrian at time $t$ onto the $(p,q)$-plane is given by
$p(-q+ct)=T_{0}+(T_{1}-T_{0})t.$
It follows that each of this family of slow thermodynamic paths passes through the point given by (31):
$(p_{*},q_{*})=\Big{(}\frac{T_{1}-T_{0}}{c},-\frac{cT_{0}}{T_{1}-T_{0}}\Big{)}.$
In addition, the negative free energy increases during this process:
$z(t)=-G(t)=T(t)\ln v(t)=(T+(T_{1}-T_{0})t)\ln v,\quad v\gg 1.$
Thus, the individual trajectory of the macroscopic state $(z_{0},p_{*},q_{*})$ is a Reeb chord. Note that the Legendrian isotopy described above is non-negative, and hence represents a legitimate thermodynamic process, only if we consider a small neighborhood of the point $(T_{0}\ln q_{*},p_{*},q_{*})$ on the initial Legendrian.
6.2 A temperature and magnetic field jump with constant magnetization
We start with a discussion based on [12]. Consider the Curie-Weiss magnet in the thermodynamic limit, see Example 3.3. As the external magnetic field $H$ is an intensive variable, and the magnetization $M$ is an extensive one, we use the notation $p=M\in{\mathbb{R}}$ and $q=H\in{\mathbb{R}}$. Recall (cf. also Remark 6.1) that the total magnetic field is given by $H+H_{back}$, where $H_{back}$ is the background magnetic field, playing a role of the parameter. Suppose that the temperature of the Curie-Weiss magnet jumped from $T_{0}$ to $T_{1}>T_{0}$, and that the background magnetic field jumped from 00 to $c\in{\mathbb{R}}$. The reduced Legendrian equilibrium submanifold $\Lambda(T,H_{back})$ is defined by the first two equations in (26). Let us set $\Lambda_{0}=\Lambda(T_{0},0)$, and let $\Lambda_{1}=\Lambda(T_{1},c)$.
We claim that there is a unique chord Reeb chord joining $\Lambda_{0}$ and $\Lambda_{1}$. Indeed, it follows from the second equation in (26) that such a chord projects to the point $(p,q)$ if and only if
$q+c+bp=T_{1}\tanh^{-1}p,~{}~{}q+bp=T_{0}\tanh^{-1}p\;,$ (34)
which gives
$p=\tanh\frac{c}{T_{1}-T_{0}}\;.$
Substituting back into either of the two equations in (34), we get
$q=\frac{cT_{0}}{T_{1}-T_{0}}-b\tanh\frac{c}{T_{1}-T_{0}}\;.$ (35)
Moreover, the free energy
$z=T\ln\Big{(}2\cosh\frac{q+bp}{T}\Big{)}-\frac{b}{2}p^{2}$
is an increasing function of $T$ for $p$ and $q$ fixed because
$\frac{\partial{z}}{\partial{T}}$ $=\ln 2+\ln\Big{(}\cosh\frac{A}{T}\Big{)}+\frac{T}{\cosh(A/T)}\sinh\Big{(}\frac{A}{T}\Big{)}\Big{(}-\frac{A}{T^{2}}\Big{)}$ (36)
$=\ln 2+\ln\cosh(A/T)-\frac{A}{T}\tanh\Big{(}\frac{A}{T}\Big{)}=\ln\Big{(}\frac{e^{A/T}+e^{-A/T}}{e^{(A/T)\tanh(A/T)}}\Big{)}>0.$
Here, we have set $A=b+pq$. Thus, the chord goes in the direction of the Reeb field $\partial/\partial z$. The claim follows.
As discussed in [12], the chords correspond to ultrafast relaxation processes in the thermodynamic limit under an additional assumption that they connect stable (as opposed to unstable or metastable) equilibria of the Curie-Weiss model, (see Remark 4.1 above). This happens, for instance, if $bT_{1}<1$, or if $c>0$ and for $q$ given by (35) both $q$ and $q+c$ have the same sign.
Keeping the previous notation, we have the following result.
Theorem 6.4.
Let $g_{t}$, $t\in[0,1]$ be a compactly supported contact isotopy of $J^{1}{\mathbb{R}}$ such that
$g_{t}(\Lambda_{1})\cap\Lambda_{0}=\emptyset\;\;\;\forall t\in[0,1]\;.$ (37)
Then there exists the Reeb chord starting at $\Lambda_{0}$ and ending at $g_{1}(\Lambda_{1})$, i.e. an instantaneous relaxation process corresponding to the above jump.
Proof.
The Legendrian $\Lambda(T,H_{back})$ can be equivalently represented as
$z=\phi_{T}(q+H_{back}+bp)-\frac{bp^{2}}{2},~{}~{}p=\phi_{T}^{\prime}(q+H_{back}+bp).$ (38)
Here, we have set
$\phi_{T}(q)=T\ln\Big{(}2\cosh\Big{(}\frac{q}{T}\Big{)}\Big{)}.$ (39)
We make a change of variables as in [11]
$Q=q+bp,\;P=p-\phi^{\prime}_{T_{0}}(Q),\;Z=z-\phi_{T_{0}}(Q)+\frac{bp^{2}}{2}\;,$ (40)
which preserves the Gibbs form:
$dZ-PdQ$ $=dz-\phi_{T_{0}}^{\prime}(Q)dq-b\phi_{T_{0}}^{\prime}(Q)dp+bpdp-(p-\phi^{\prime}_{T_{0}}(Q))(dq+bdp)$ (41)
$=dz+(-\phi_{T_{0}}^{\prime}(Q)-p+\phi_{T_{0}}^{\prime}(Q))dq+(-b\phi_{T_{0}}^{\prime}(Q)+bp-bp+\phi^{\prime}_{T_{0}}(Q))dp$
$=dz-pdq.$
Hence, it also preserves the Reeb chords. It follows from (38) and (40) that under this change of variables, the Legendrian submanifold $\Lambda_{0}=\Lambda(T_{0},0)$ becomes the zero section
$L=\{Z=0,~{}P=0\}.$
The Legendrian $\Lambda_{1}=\Lambda(T_{1},c)$ after the change of variables becomes
$K:=\{Z=\psi(Q):=\phi_{T_{1}}(Q+c)-\phi_{T_{0}}(Q),\;P=\psi^{\prime}(Q)\}\;.$ (42)
An elementary analysis shows that $\psi$ is asymptotic to the lines $\{z=\pm c\}$ as $Q\to\pm\infty$, and has its unique maximum at $Q_{*}=cT_{0}/(T_{1}-T_{0})$, compare with the previous section and see Fig. 4. This maximum corresponds to a Reeb chord from the zero section $L$ to $K$.
Figure 4: $b=1,\beta_{0}=0.5,\beta_{1}=0.3,c=1$
Let now $g_{t}$, $t\in[0,1]$ be a compactly supported contact isotopy of $J^{1}{\mathbb{R}}$ satisfying (37). Fix sufficiently large positive numbers $R>r>0$, let $A_{R}$ be the cube $[-R,R]^{3}\subset{\mathbb{R}}^{3}=J^{1}{\mathbb{R}}$ and denote by $S_{R}$ the circle ${\mathbb{R}}/R{\mathbb{Z}}$ in the $Q$-variable. For a Legendrian submanifold $Y$ which coincides with $K$ outside $A_{r}$, we denote by $\widehat{Y}$ the Legendrian submanifold of $J^{1}S_{R}$ obtained by the following procedure (see Fig. 5). First, interpolate between $Y$ and the Legendrian submanifold
$L_{c}=\{z=-c,p=0\}$
inside the interior of $A_{R}$ without creating new chords from the zero section $L$ to the resulting submanifold $\widehat{Y}$. Then, glue together the points of this submanifold corresponding to $Q=-R$ and $Q=R$. This is possible simply because $L_{c}$ does not depend on $Q$.
Figure 5: Interpolation and gluing in $(Z,Q)$-plane
Observe that there is a unique non-degenerate chord from $\widehat{L}$ to $\widehat{K}$ corresponding to $Q_{*}$, and that these submanifolds are explicitly Legendrian isotopic by changing $T$ from $T_{0}$ to $T_{1}$. Thus, by a result from [10], the submaniflds $\widehat{L}$ to $\widehat{K}$ are interlinked (see [10, 11] for the definition of interlinking). In particular, if $g_{t}$, $t\in[0,1]$, is a contact isotopy of ${\mathbb{R}}^{3}$ supported in $Q_{r}$ such that $g_{t}(K)$ does not touch $L$ for all $t$, it descends to $J^{1}S_{R}$, and hence $\widehat{g_{1}(K)}$ and $\widehat{L}$ are interlinked. Since the interpolation above does not create new chords, we get a chord between $g_{1}(K)$ and $L$. Returning to the original coordinates $(z,p,q)$ we get the statement of the theorem. ∎

### v3

@@
 One natural question that can be addressed by the tools of contact topology is the existence of an ultrafast processes that would connect the equilibria of two thermodynamic systems.
-As we discuss in Theorem 5.2 below, if two equilibrium thermodynamic Legendrians are related by a slow temperature non-decreasing global process, there is an ultrafast process which starts at an equilibrium state of the first system and ends at an equilibrium state of the second system.
+As we discuss in Theorem 5.2 below, if two equilibrium thermodynamic Legendrians are related by a slow temperature non-decreasing global process, there is a free-energy non-increasing ultrafast process which starts at an equilibrium state of the first system and ends at an equilibrium state of the second system.
 Remark 1.1.
-It should be mentioned that when the process is not quasi-static (that is, if we cannot assume that at each time moment the system is in the equilibrium), the temperature of the system is not well defined from the physical point of view.
+It should be mentioned that when the process is not quasi-static (that is, if we cannot assume that at each moment of time the system is in the equilibrium), the temperature of the system is not well defined from the physical point of view.
 We always assume that the system exchanges heat with a thermostat (or reservoir) which has a well defined temperature, and for a system not in an equilibrium, we will consider the temperature of the reservoir at the given time.
@@
 The definitions of the slow, fast and ultrafast processes are discussed in Section 4.
-Section 5 uses some non-trivial results in symplectitc topology to obtain certain consequences of the existence of slow global processes connecting the equilibrium Legendrian submanifolds for two given thermodynamic systems.
+Section 5 uses some non-trivial results in symplectic topology to obtain certain consequences of the existence of slow global processes connecting the equilibrium Legendrian submanifolds for two given thermodynamic systems.
 Finally, in Section 6, we discuss the ultrafast processes in the examples of the ideal gas and Stirling engine, and the Curie-Weiss magnet.
@@
 Acknowledgment.
-We thank Shin-itiro Goto for very useful comments on the manuscript.
-Preliminary results of this paper were presented by L.P.
-at the Symplectic Geometry Seminar in IAS, Princeton.
+We thank Bernard Derrida, Leonid Levitov, Shin-itiro Goto, and Nikita Nekrasov for very useful comments.
+Preliminary results of this paper were presented by LP at the Symplectic Geometry Seminar in IAS, Princeton.
 The recording of this talk, called “Contact topology meets thermodynamics”, is available on YouTube.
+The paper was written during LP’s sabbatical at the University of Chicago.
+LP thanks the Department of Mathematics for an exceptional research atmosphere.
 ME was partially supported by the Israel Science Foundation grant 2033/23, LP was partially supported by the Israel Science Foundation grant 1102/20, and LR was partially supported by NSF grant DMS-2205497 and by ONR grant N00014-22-1-2174.
@@
 2.1 Microscopic and macroscopic thermodynamic states
-[…]
+A microscopic state of a thermodynamic system is a point on a manifold $M$, equipped with a measure $\mu$.
+We let $\mathcal{P}$ be the collection of smooth positive densities $\rho>0$ on $M$ so that $\int_{M}\rho d\mu=1$, and refer to an element $\rho\in\mathcal{P}$ as a macroscopic state.
+Its entropy is
 $S(\rho)=-\int_{M}\rho\ln\rho~d\mu,~~{\rho\in{\mathcal{P}}}.$ (1)
@@
 The free energy of the system is
-[…]
+$G(T,q,\rho)=-TS(\rho)+\int_{M}H(q,m)\rho(m)d\mu(m),~~q\in{\mathbb{R}}^{n},~\rho\in{\mathcal{P}}.$ (2)
 Here, $q\in{\mathbb{R}}^{n}$ is an exterior physical parameter and $T>0$ is the temperature of the system (see Remark 1.1). The Hamiltonian $H(q,m)$ often appears in the form
-$H(q,m)=V_{int}(m)+V_{ext}(q,m),$ (3)
-where $V_{int}(m)$ is the microscopic internal energy, and $V_{ext}(q,m)$ is related to an external influence on the system, as indicated by its dependence on the exterior parameter $q\in{\mathbb{R}}^{n}$.
+$H(q,m)=V_{\rm int}(m)+V_{\rm ext}(q,m),$ (3)
+where $V_{\rm int}(m)$ is the microscopic internal energy, and $V_{\rm ext}(q,m)$ is related to an external influence on the system, as indicated by its dependence on the exterior parameter $q\in{\mathbb{R}}^{n}$.
 A special case of interest is when $V(q,m)$ is linear in $q$:
-$V_{ext}(q,m)=(q\cdot\overline{V}(m)).$ (4)
-Here, ()($\cdot$)( ⋅ ) denotes the standard inner product on ${\mathbb{R}}^{n}$, and $\overline{V}$ is a map from $M$ to ${\mathbb{R}}^{n}$.
+$V_{\rm ext}(q,m)=(q\cdot\overline{V}(m)).$ (4)
+Here, $(\cdot)$ denotes the standard inner product on ${\mathbb{R}}^{n}$, and $\overline{V}$ is a map from $M$ to ${\mathbb{R}}^{n}$.
 For every $T>0$, and $q\in{\mathbb{R}}^{n}$ fixed, the functional $G(q,\cdot)$ is strictly convex on $\mathcal{P}$.
 A direct computation shows that its unique critical point, known as the Gibbs distribution, or an equilibrium distribution, is
-[…]
+${\rho_{G}:=\hbox{argmin}_{\rho\in{\mathcal{P}}}G(T,q,\rho)}=\frac{e^{-\beta H(q,m)}}{\mathcal{Z}(q)},~~\mathcal{Z}(q)=\int_{M}e^{-\beta H(q,m)}d\mu(m).$ (5)
 Note that the entropy of a macroscopic state $\rho\in{\mathcal{P}}$ can be written as
@@
 Similarly, for given a macroscopic state $\rho\in{\mathcal{P}}$ the generalized pressures [3] are defined as
-[…]
+$p_{j}{(q,\rho)}=-\frac{\partial G}{\partial q_{j}}{(T,q,\rho)}=-\int_{M}\frac{\partial H}{\partial q_{i}}{(q,m)}\rho{(m)}d\mu({m})\;,\;j=1,\dots,n\;.$ (7)
 Note that $p_{i}$ do not depend on $T$.
@@
 In the special case of the linear Hamiltonians as in (3)-(4), the pressures have the form
-[…]
-and free energy in (2) is
+$p_{j}(\rho)=-\int_{M}\bar{V}_{j}(m)\rho(m)d\mu(m),~~j=1,\dots,n,$
+and the free energy in (2) is
 $G(T,q,\rho)$ $=U(\rho)-TS(\rho)-\sum_{j=1}^{n}p_{j}(\rho)q_{j}.~~q\in{\mathbb{R}}^{n},~\rho\in{\mathcal{P}},$ (8)
 where $U(\rho)$, called the macroscopic internal energy, is defined by
-[…]
+$U(\rho)=\int_{M}V_{\rm int}(m)\rho(m)d\mu(m).$ (9)
 2.2 A geometric interpretation of the equilibria and the extended thermodynamic phase space
 These basic notions can be represented geometrically as follows.
-Given a macroscopic state $\rho\in{\mathcal{P}}$, we consider the point $(-G(T,q,\rho),S(\rho),T,p(q,\rho),q)$ as an element of ${\mathbb{R}}^{2n+3}$ with the coordinates $z=-G$, $S,T$, and $(p_{j},q_{j})$, $j=1,\ldots,n$.
+Given a macroscopic state $\rho\in{\mathcal{P}}$, we consider the point $(-G(T,q,\rho),S(\rho),T,p(q,\rho),q)$ as an element of ${\mathbb{R}}^{2n+3}$ with the coordinates z=Gz=-G, $S,T$, and $(p_{j},q_{j})$, $j=1,\ldots,n$.
 This space is called the extended thermodynamic phase space and is denoted by $\widehat{\mathcal{T}}$.
@@
 $\widehat{\lambda}=dz-SdT-\sum_{j=1}^{n}p_{j}dq_{j},$ (10)
-called the Gibbs 1-form.
-It is a contact form: the maximal dimension of an integral submanifold $\Lambda$ of the distribution $\text{ker}(\lambda)$ equals $n+1$.
-Recall that $\Lambda$ is called integral if it is tangent to the distribution, that is, $\lambda$ vanishes on $T\Lambda$.
-The integral submanifolds of the maximal dimension $n+1$ are called Legendrian.
+[…]
 In our setting, a standard example of a Legendrian submanifold is the 1-jet of a smooth function $f:{\mathbb{R}}^{n+1}\to{\mathbb{R}}$,
-$\Lambda_{f}=\Big{\{}(z,S,T,p,q,z)\in{{\mathbb{J}}^{1}{\mathbb{R}}^{n+1}}\Big{|}~{}z=f(T,q),~{}S=\frac{\partial{f}}{\partial{T}}(T,q),~{}p_{j}=\frac{\partial f}{\partial q_{j}}(T,q),j=1,\ldots,n\Big{\}}\;.$ (11)
+$\widehat{\Lambda}_{f}=\Big\{(z,S,T,p,q,z)\in{{\mathbb{J}}^{1}{\mathbb{R}}^{n+1}}\Big|~z=f(T,q),~S=\frac{\partial{f}}{\partial{T}}(T,q),~p_{j}=\frac{\partial f}{\partial q_{j}}(T,q),j=1,\ldots,n\Big\}\;.$ (11)
 More generally, given a smooth function
-$\Psi:{\mathbb{R}}^{n+1}\times E\to{\mathbb{R}}\;,$
-where $E$ is a space of auxiliary variables $\xi$, the set $\Lambda_{\Psi}$ of all $(z,S,T,p,q)\in{\mathbb{J}}^{1}{\mathbb{R}}^{n+1}$ such that there exists $\xi\in E$ so that
-$z=\Psi(T,q,\xi),~{}\frac{\partial\Psi}{\partial\xi}(T,q,\xi)=0,~{}S=\frac{\partial\Psi}{\partial T}(T,q,\xi),~{}p=\frac{\partial\Psi}{\partial q}(T,q,\xi),$ (12)
-is a (possibly singular) Legendrian submanifold for a generic $\Psi$. The function $\Psi$ is called a generating function of $\Lambda_{\Psi}$, and the variables $\xi$ are called the ghost variables.
+$\Psi:{\mathbb{R}}^{n+1}\times{\mathcal{A}}\to{\mathbb{R}}\;,$
+where $\mathcal{A}$ is a linear space of auxiliary variables $\xi$, the set $\Lambda_{\Psi}$ of all $(z,S,T,p,q)\in{\mathbb{J}}^{1}{\mathbb{R}}^{n+1}$ such that there exists $\xi\in\mathcal{A}$ so that
+$\widehat{\Lambda}_{\Psi}=$ $\Big\{\hbox{$(z,S,T,p,q):\exists\xi\in\mathcal{A}$ s.t }z=\Psi(T,q,\xi),~\frac{\partial\Psi}{\partial\xi}(T,q,\xi)=0,$~ (12)
+$S=\frac{\partial\Psi}{\partial T}(T,q,\xi),~p=\frac{\partial\Psi}{\partial q}(T,q,\xi)\Big\},$
+is a (possibly singular) Legendrian submanifold for a generic $\Psi$.
+The function $\Psi$ is called a generating function of $\widehat{\Lambda}_{\Psi}$, and the variables $\xi$ are called the ghost variables.
 A key observation is that (5)-(7) exactly mean that a Gibbs measure $\rho_{G}(T,q,\cdot)$ of a thermodynamic system corresponds to a point on the Legendrian submanifold $\Lambda$ of the form (12) with the elements $\rho\in{\mathcal{P}}$ being the ghost variables $\xi$.
-The generating function of $\Lambda$ is the (minus) free energy $-\Phi$ (we learned this observation from [23]).
+The generating function of $\Lambda$ is the (minus) free energy $\Phi$ (we learned this observation from [27]).
 In other words, the equilibria of a thermodynamic system form a Legendrian submanifold $\widehat{\Lambda}$ in $\widehat{\mathcal{T}}$, as we vary the external parameters $T$ and $q$, that themselves are coordinates on $\widehat{\mathcal{T}}$.
@@
 This is an $n+2+k$ dimensional sub-space.
-The reduced thermodynamic phase space $\mathcal{T}$ is obtained by the projection of $h$ along the directions corresponding to the variables $S$, $p_{i},i\in I$ and $q_{e},e\in E$, and is $12k+$ 1 dimensional.
+The reduced thermodynamic phase space $\mathcal{T}$ is obtained by the projection of $h$ along the directions corresponding to the variables $S$, $p_{i},i\in I$ and $q_{e},e\in E$, and is $2k+1$ dimensional.
+Remark 2.1.
+Note that our notions of extended and reduced thermodynamic phase spaces and of reduction differ from the notions with the same names appearing in some other papers on geometric approaches to thermodynamics - see e.g.
+[36].
 The reduced thermodynamic phase space is naturally identified with $J^{1}{\mathbb{R}}^{k}$. Consider now the projection $J^{1}{\mathbb{R}}^{n+1}\to J^{1}{\mathbb{R}}^{k}$,
@@
 A path $\gamma(t)\subset J^{1}{\mathbb{R}}^{n+1}$ is called non-negative if $\widehat{\lambda}(\dot{\gamma}(t))\geq 0$ for every $t$.
+Let us emphasize that the definition does not depend on the choice of the specific contact form $\widehat{\lambda}$ but only on the contact distribution and its co-orientation.
 To derive the aforementioned constraint, let $(T(t),q(t)),\rho(t))$ be a smooth path and $\dot{\gamma}(t)$ and $\dot{\rho}(t)$ be, respectively, the tangent vectors to the corresponding thermodynamic path $\gamma(t)$ in (15) and to $\rho(t)\in{\mathcal{P}}$.
@@
 This is because we always have $S(t)\geq 0$ by the definition (1) of the entropy.
-Similarly, in the ideal gas example considered in Section 6.1 below, the volume $V$ and the negative pressure $-P$ are pair-wise conjugate extensive and intensive variables, respectively, and the volume is, of course, always non-negative.
+Similarly, in the ideal gas example considered in Section 6.1 below, the volume $V$ and the negative pressure P-P are pair-wise conjugate extensive and intensive variables, respectively, and the volume is, of course, always non-negative.
 Thus, if we reduce the pressure only, the assumption above reads that the pressure is non-increasing.
@@
 As $p_{e}=0$ for $e\in E$, we have
-$0\leq\hat{\lambda}\Big{(}\frac{d\widehat{\gamma}}{dt}(t)\Big{)}$ $=\dot{z}(t)-S(t)\dot{T}(t)-\sum_{i=j}^{n}p_{j}(t)\dot{q}_{j}(t)$ (23)
+$0\leq\hat{\lambda}\Big(\frac{d\widehat{\gamma}}{dt}(t)\Big)$ $=\dot{z}(t)-S(t)\dot{T}(t)-\sum_{j=1}^{n}p_{j}(t)\dot{q}_{j}(t)$ (23)
 $=\dot{z}(t)-S(t)\dot{T}(t)-\sum_{j=1}^{k}p_{j}(t)\dot{q}_{j}(t)-\sum_{j=k+1}^{l}p_{j}(t)\dot{q}_{j}(t).$
@@
 4 Classification of thermodynamic processes
-As we have already stated in the introduction, the main theme of this note is modeling of the non-equilibrium thermodynamic processes.
-We distinguish between three types of irreversible processes in the reduced (at least with respect to temperature) thermodynamic phase space.
-We adopt Assumption 3.2.
+For our discussion on non-equilibrium thermodynamics, we have to deal with systems depending on an external parameter from a subset of a Euclidean space.
+For instance, we always reduce the entropy and the temperature, considering the latter to be a parameter.
+As far as the other external parameters $\alpha$ are concerned, we assume that the Hamiltonian depends on $\alpha$ in the following way:
+$(q,m,\alpha)\mapsto V^{\prime}_{\rm int}(m,\alpha)+q\cdot\overline{V}(m)\;,$
+i.e., that only the internal energy $V^{\prime}_{\rm int}$ depends on $\alpha$.
+Here is an important example.
+It is convenient to introduce an auxiliary external parameter to describe a jump of an extensive variable.
+Take a Hamiltonian $H$ of the form (3)-(4) and write a new Hamiltonian, depending on $\alpha$, as
+$H(q+\alpha,m)=V^{\prime}_{\rm int}(m,\alpha)+q\cdot\overline{V}(m)=(V_{\rm int}(m)+\alpha\cdot\overline{V}(m))+q\cdot\overline{V}(m),$
+where $V_{\rm int}(m)$ is the internal energy term in the original $H$ and $V^{\prime}_{\rm int}(m,\alpha):=V_{\rm int}(m)+\alpha\cdot\overline{V}(m)$ is the new internal energy depending on $\alpha$.
+With this notation, the jump $q\to q+c$ can be described as the jump of $\alpha$ from 0 to $c$.
+We refer to Remark 6.2 for further discussion.
+Sometimes, we consider genuine external parameters, such as the exchange parameter of the Curie-Weiss magnet, see Section 6.3 below.
+We introduce several types of non-equilibrium processes.
+They differ by their duration - slow (quasi-static), fast, and ultrafast (instantaneous), and by their “spread” in the thermodynamic phase space - local and global.
 I.
-Slow (quasi-static) global processes: They are given by non-negative families (or paths) of Legendrian (equilibrium) submanifolds.
+Slow (quasi-static) processes.
+They are caused by a slow change of some parameters of the system.
+The change is so slow that, at each moment in time, the system is in equilibrium.
+We adopt Assumption 3.2, meaning that the decrement of the free energy due to the change of the reduced intensive variables is non-negative.
+In particular, when the temperature and the entropy are the only reduced variables, we assume that the temperature is non-decreasing.
+Thus, the slow processes are given by non-negative families (or paths) of Legendrian (equilibrium) submanifolds.
 Mathematically, this means that we have an initial Legendrian submanifold $\Lambda\subset\mathcal{T}$ and a map
 $\Gamma:\Lambda\times[0,\tau]\to\mathcal{T},\;\Gamma(x,0)=x\;\;\forall x\in\Lambda,$
-such that for each $t$, the image of $\Gamma(\cdot,t)$ is a Legendrian submanifold – such a map $\Gamma$ is called a (parameterized) Legendrian isotopy, see Section 5 below222 As we already stated in the introduction, by a global process we mean a process defined at all macroscopic equilibrium states of the system.
+such that for each $t$, the image of $\Gamma(\cdot,t)$ is a Legendrian submanifold – such a map $\Gamma$ is called a (parameterized) Legendrian isotopy, see Section 5 below.
 In the geometric language, this means that the map $\Gamma$ describing this process is defined on the direct product of the whole equilibrium Legendrian submanifold $\Lambda$ – and not only of its part – with the time interval.
-All Legendrian submanifolds that we consider are assumed to be closed subsets of the thermodynamic phase space..
-Furthermore, we assume that each individual path $t\mapsto\Gamma(x,t)$ (for a fixed $x$) is non-negative, that is, a thermodynamic path satisfying (19).
-(Such a Legendrian isotopy is called non-negative, see Section 5 below).
-Physically, this means that the change is so slow that, at each moment in time, the system is in equilibrium.
+We assume that each individual path $t\mapsto\Gamma(x,t)$ (for a fixed $x$) is non-negative, that is, a thermodynamic path satisfying (24).
+Such a Legendrian isotopy is called non-negative, see Section 5 below.
+A thermodynamic path $t\mapsto\Gamma(x,t)$ appearing in a slow (quasi-static) process $\Gamma$ will be called a slow thermodynamic path.
+We distinguish between two kinds of slow processes.
+We call a slow process global if the isotopy $\Gamma$ is a proper embedding (i.e., preimage of every compact subset is compact).
+[^2] As we already stated in the introduction, by a global process we mean a process defined at all macroscopic equilibrium states of the system.
 An example is provided by the Curie-Weiss magnet whose temperature is slowly increasing, yielding a path of submanifolds $\Lambda(T,0)$ from Example 3.3.
-A thermodynamic path $t\mapsto\Gamma(x,t)$ appearing in a slow (quasi-static) global process $\Gamma$ will be called a slow thermodynamic path.
+Otherwise we call a slow process local.
+Next, we pass to a discussion of fast and ultrafast (instantaneous) processes.
+They are triggered by an abrupt change of the external parameters of the system, say $\alpha_{0}\to\alpha_{1}$.
+Assume that the initial system was in the equlibrium with the Gibbs distribution $\rho_{G,\alpha_{0}}$.
+The corresponding point in the thermodynamic phase space is $(z^{(0)},p^{(0)},q)$ with $z^{(0)}=-G(q,\rho_{G,\alpha_{0}},\alpha_{0})$ and
+$p^{(0)}=-\int_{M}\overline{V}(m)\rho_{G,\alpha_{0}}d\mu\;.$ (27)
+What happens to the system next can be described as follows.
+Two stage scenario:
+Stage 1: Jump. The macroscopic parameters of the system after the jump are $(z^{(1)},p^{(1)},q)$ with $z^{(1)}=-G(q,\rho_{G,\alpha_{0}},\alpha_{1})$ and
+$p^{(1)}=-\int_{M}\overline{V}(m)\rho_{G,\alpha_{0}}d\mu=p^{(0)},$ (28)
+cf.
+(27).
+It follows that only the free energy changes.
+Note that the system remains in the macroscopic state $\rho_{G,\alpha_{0}}$, while the microscopic internal energy instantaneously changes from $V_{\rm int}(m,\alpha_{0})$ to $V_{\rm int}(m,\alpha_{1})$.
+Stage 2: The Fokker-Planck evolution. It belongs to the class of fast processes which we are going to discuss it right now.
 II.
@@
 The corresponding path $\gamma(t)=(z(t),p(t),q)$, with $z=-G(T,q,\rho)$, converges to the equilibrium submanifold corresponding to the terminal temperature $T_{\infty}$ of the process.
-We claim that $\gamma(t)$ necessarily non-negative. Indeed, recalling (6) and since $q$ is kept fixed, we get
+We claim that $\gamma(t)$ is necessarily non-negative. Indeed, recalling (6) and since $q$ is kept fixed, we get
 $\lambda(\dot{\gamma})=\dot{z}=d_{\rho}G(\nabla_{\rho}G)+S\dot{T}\geq 0\;,$ (30)
 and the claim follows.
-An example is given by the thermodynamic path corresponding to the transition from the metastable to the stable equilibrium of the Curie-Weiss magnet, see [12] for a discussion of metastability and further references.
+One can slightly generalize the notion of the fast process by allowing not only the temperature, but also the external parameter $\alpha(t)$ to move with time.
+Let us also emphasize that here we do not assume that at time moment 0 the system is in the equilibrium.
+For instance, at Stage 2 above the evolution starts after the jump, and hence in general out of equilibrium.
+Another example is given by the thermodynamic path corresponding to the transition from the metastable to the stable equilibrium of the Curie-Weiss magnet, see [13] for a discussion of metastability and further references.
 III.
-Ultrafast (instantaneous) processes: They correspond to Reeb chords, i.e., to pieces of trajectories of the Reeb vector field $\partial/\partial z$ connecting the initial and the terminal equilibrium Legendrian submanifolds; see [12] for a detailed discussion as well as Section 6 below.
-Both variables $(p,q)$ remain constant along such a process, while $z$ is increasing.
+Ultrafast (instantaneous) processes: In the thermodynamic limit of certain mean field systems the two stage scenario above may end already at Stage 1.
+This happens when the limiting and the terminal Gibbs distributions coincide, so the fast process forming Stage 2 is void: the whole process consists of a jump only.
+Assume furthermore that these Gibbs distributions are given by a delta function on the space of microstates:
+$\rho_{G,\alpha_{0}}=\rho_{G,\alpha_{1}}=\delta(m-m_{*})\;.$ (31)
+The latter assumption can be verified for the Curie-Weiss magnet, and for the ideal gas.
+It holds true for a more general class of models, essentially due to a version of the large deviation principle.
+This will be discussed in a forthcoming paper.
+By (27) and (28) these assumptions would yield $p:=p^{(0)}=p^{(1)}=\overline{V}(m^{*})$.
+Note now that the points $(p,q,z^{(0)})$ and $(p,q,z^{(1)})$ are endpoints of the Reeb chord (i.e., a segment parallel to the $z$-axes) connecting the equilibrium Legendrians.
+Remark 4.1.
+The free energy can change either way in the course of an ultrafast process.
+If it decreases, i.e., $z_{1}>z_{0}$, there exists a local slow process consisting of germs of Legendrians “sliding” along the chord.
+This happens, for instance, with the isochoric piece of the Stirling engine.
+It can be realized by a slow local isochoric process in the quasitatic limit (i.e., when the duration of the cycle tends to infinity), cf.
+[34].
 Remark 4.2.
@@
 We do not address the issue of the stability of equilibria in the context of Reeb chords, and it would be interesting to understand if contact topology can be helpful here.
-For simplicity, we stick to the terminology in III.
 Remark 4.3.
@@
 This is illustrated in Example 6.4 below.
+Let us compare our description of thermodynamic processes with some earlier literature.
+Some authors propose to model non-equilibrium processes by contact Hamiltonian flows, see for instance [5, 17, 12, 18, 19, 20].
+This approach proved successful in several interesting cases, illustrating in particular a thesis from Prigogine’s 1977 Nobel lecture [32]—that thermodynamic potentials act as Lyapunov functions near equilibrium.
+At the same time it has limitations in the presence of phase transitions [12].
+In particular, one can prove non-existence (even locally!) of time-independent contact Hamiltonian that models some natural relaxation processes [20].
+Interestingly enough, positive Legendrian isotopies, proposed as models of slow thermodynamic processes in the present paper can be realized non-uniquely by (in general) non-autonomous ambient contact Hamiltonian flows.
+This follows from the isotopy extension theorem in contact topology [16].
+Highlighting the relevance of positive Legendrian isotopies, and in particular the link with the partial order on the space of Legendrians, is one of the main findings of the present paper.
+For two different contact viewpoints at the Fokker-Planck evolution (cf.
+the second stage of the scenario above) we refer to [13] and [19].
+Let us mention also that the GENERIC formalism [21, 22] provides a simultaneous description of reversible and irreversible thermodynamic processes, including the Fokker-Planck evolution.
+The interpretation of the ultrafast processes as Reeb chords between the initial and the terminal Lagrangians appeared first in [13] and is further developed in the present paper.
+One can analyze it through the lens of the correspondence between the quantum mechanics and thermodynamics (we thank N.Nekrasov for this suggestion).
+Recall that this correspondence is given by the passage to imaginary time $i\hbar/T$ which is proportional to the inverse temperature, see e.g.
+[14, Section 10-2] and [33, formula(2.285)].
+Assume for simplicity that the Lagrangian projections $L_{0}$ and $L_{1}$ of the initial and the terminal Legendrians $\Lambda_{0}$ and $\Lambda_{1}$, respectively, are graphical of the form
+$L_{j}=\Big\{p=\frac{\partial f_{j}}{\partial q}\Big\}\;,\;j=0,1\;,q\in{\mathbb{R}}^{n},$
+and that they intersect transversely in a finite number of points.
+On the quantum side $L_{j}$ defines the space of Lagrangian (or WKB) states of the form $a_{j}(q)e^{i\frac{f_{j}(q)}{\hbar}}\;,$ with $\int|a_{j}|^{2}dq=1$, see e.g.
+[2].
+The complex-valued functions $a_{j}$ are assumed to be smooth and compactly supported.
+We will denote such a state $\xi_{j}(a_{j},\hbar)$.
+The transition probability between $\xi_{0}(a_{0},\hbar)$ and $\xi_{1}(a_{1},\hbar)$ is given by
+$p(a_{0},a_{1},\hbar):=|\langle\xi_{0}(a_{0},\hbar),\xi_{1}(a_{1},\hbar)\rangle|^{2}\;.$
+Here $\langle\cdot,\cdot\rangle$ is the $L_{2}$-scalar product with respect to the Euclidean volume. The stationary phase formula (see e.g. [23, Section 15.6.2]) yields
+$p(a_{0},a_{1},\hbar)=\hbar^{n/2}\int a_{0}\overline{a_{1}}\exp{\left(\frac{i}{\hbar}(f_{0}-f_{1})\right)}d\mu+O(\hbar^{1+n/2})\;,$ (32)
+where $\mu$ is a complex-valued measure.
+Crucially, $\mu$ supported in a finite set $\mathcal{Q}$ of the $q$-coordinates of the intersection points from $L_{0}\cap L_{1}$.
+Note that (tautologically!) this set is in a one to one correspondence with the set of chords connecting $\Lambda_{0}$ and $\Lambda_{1}$.
+In fact, the points of $\mathcal{Q}$ are exactly the special values of intensive variables corresponding to ultrafast processes.
+It would be interesting to explore this analogy further.
 5 Slow global processes and partial order on Legendrians
@@
 We say that a Legendrian isotopy is compactly supported if it admits a compactly supported parameterization.
-Let $[\Lambda]$ be the family of all Legendrians in $(\Sigma,\xi)$ that can be obtained from a given properly embedded Legendrian $\Lambda$ by a compactly supported Legendrian isotopy.
-We say that a compactly supported Legendrian isotopy $\Lambda_{t}$, $t\in[0,1]$, of Legendrians in $[\Lambda]$ is non-negative if it admits a parameterization with a vector field $v_{t}$ satisfying $\alpha(v_{t})\geq 0$ (note that if this holds for one parameterization, then it holds for all the others).
+Let $\Lambda]$ be the family of all Legendrians in $(\Sigma,\xi)$ that can be obtained from a given properly embedded Legendrian $\Lambda$ by a compactly supported Legendrian isotopy.
+We say that a compactly supported Legendrian isotopy $\Lambda_{t}$, $t\in[0,1]$, of Legendrians in $\Lambda]$ is non-negative if it admits a parameterization with a vector field $v_{t}$ satisfying $\alpha(v_{t})\geq 0$ (note that if this holds for one parameterization, then it holds for all the others).
 We write $\Lambda_{0}\preceq\Lambda_{1}$ if there exists a non-negative Legendrian isotopy from $\Lambda_{0}$ to $\Lambda_{1}$.
@@
 After stating them, together with their thermodynamic interpretations, we outline their proofs and provide relevant references.
-The class $[\Lambda]$ is called orderable if the binary relation $\preceq$ is a partial order.
+The class $\Lambda]$ is called orderable if the binary relation $\preceq$ is a partial order.
 A Reeb chord between two Legendrians is called non-trivial if it has positive length (as opposed to a point).
@@
 Theorem 5.1.
-For any (open or closed) smooth connected manifold $X$, the class $[0_{X}]$ of the zero section $0_{X}$ in $J^{1}X=T^{*}X\times{\mathbb{R}}$ is orderable.
+For any (open or closed) smooth connected manifold $X$, the class $0_{X}]$ of the zero section $0_{X}$ in $J^{1}X=T^{*}X\times{\mathbb{R}}$ is orderable.
 If two distinct equilibrium Legendrians $\Lambda_{0},\Lambda_{1}\in[0_{X}]$ are related by the partial order, $\Lambda_{0}\preceq\Lambda_{1}$, there is no admissible global process taking $\Lambda_{1}$ to $\Lambda_{0}$.
@@
 If two distinct Legendrians $\Lambda_{0},\Lambda_{1}\in[0_{X}]$ satisfy $\Lambda_{0}\preceq\Lambda_{1}$, there exists a non-trivial Reeb chord starting on $\Lambda_{0}$ and ending on $\Lambda_{1}$.
-In particular, if two equilibrium thermodynamic Legendrians from $[0_{X}]$ are related by an admissible global process, then there is an ultrafast process which starts at an equilibrium state of the first system and ends at an equilibrium state of the second system.
+In particular, if two equilibrium thermodynamic Legendrians from $0_{X}]$ are related by an admissible global process, then there is a free-energy non-increasing ultrafast process which starts at an equilibrium state of the first system and ends at an equilibrium state of the second system.
 Theorem 5.3.
-For every two distinct Legendrians in $[0_{X}]$ there exists a non-trivial Reeb chord connecting them (in an unspecified order).
+For every two distinct Legendrians in $0_{X}]$ there exists a non-trivial Reeb chord connecting them (in an unspecified order).
 If these are the equilibrium Legendrians of two thermodynamic systems, there exists an ultrafast process that starts at an equilibrium state of one system and ends at an equilibrium state of the other.
@@
 Case I: $X$ is a closed manifold.
-In this case, Theorem 5.1 follows from [5, Proposition 5], [7, Theorem 1].
+In this case, Theorem 5.1 follows from [6, Proposition 5] and [8, Theorem 1].
 Theorem 5.2 can be deduced from the theory of generating functions and their spectral selectors [7, 8, 38] along the following lines.
 Assume without loss of generality that $\Lambda_{0}$ is the zero section.
-Let $c_{+}$ and $c_{-}$ be the spectral selectors corresponding to the fundamental class (i.e., the class of $X$) and the class of the point in the homology $H_{*}(X,\mathbb{Z}_{2})$ with the $Z_{2}$-coefficients, respectively.
+Let $c_{+}$ and $c_{-}$ be the spectral selectors corresponding to the fundamental class (i.e., the class of $X$) and the class of the point in the homology $H_{*}(X,\mathbb{Z}_{2})$ with the $\mathbb{Z}_{2}$-coefficients, respectively.
 If $\Lambda_{0}\preceq\Lambda_{1}$, then $c_{+}(\Lambda_{1})\geq 0$.
@@
 Proposition 2.4(i) in [38]) that there exists a chord of time-length $c_{+}(\lambda_{1})$ starting on $\Lambda_{0}$ and ending on $\Lambda_{1}$, as required.
-Theorem 5.3 follows from the fact that since $\Lambda_{0}\neq\Lambda_{1}$, at least one of $c_{-}(\Lambda_{1})$ or $c_{+}(\Lambda_{1})$ is distinct from 00.
-A different argument which deduces Theorems 5.2 and 5.3 from the orderability of the class $[0_{X}]$ is discovered in a recent paper by Allais and Arlove [1], see Corollary 1.5 and Theorem 4.7 therein.
-Interestingly enough, this argument works on arbitrary contact manifolds, with $[0_{X}]$ being replaced by arbitrary orderable Legendrian isotopy class of a closed Legendrian submanifold.
+Theorem 5.3 follows from the fact that since $\Lambda_{0}\neq\Lambda_{1}$, at least one of $c_{-}(\Lambda_{1})$ or $c_{+}(\Lambda_{1})$ is distinct from 0.
+A different argument which deduces Theorems 5.2 and 5.3 from the orderability of the class $0_{X}]$ is discovered in a recent paper by Allais and Arlove [1], see Corollary 1.5 and Theorem 4.7 therein.
+Interestingly enough, this argument works on arbitrary contact manifolds, with $0_{X}]$ being replaced by arbitrary orderable Legendrian isotopy class of a closed Legendrian submanifold.
 Our exposition is influenced by this paper.
@@
 On the other hand, it is unclear how to characterize or detect such pairs in general.
-Remark 5.5.
+Remark $5.$
 We refer to [28] for appearance of some other order relations in thermodynamics. It would be interesting to relate the discussion in [28] with the partial order on Legendrians considered above.
@@
 Furthermore, we assume that the gas is diluted, $v\gg 1$.
-This is a model approximating the ideal gas – see [13, Section 1.3.2] for details.
+This is a model approximating the ideal gas – see [15, Section 1.3.2] for details.
 In this model, the temperature $T$ and the pressure $P$ are the intensive variables, while the volume $v$ is the extensive variable.
@@
 If the Bolzmann and the gas constants are taken to be $1$, in an equilibrium we have
-$(P+P_{back})v=T.$ (29)
-Here, $P_{back}$ is a background pressure that we treat as a physical parameter of the system.
-If the molecules do not interact so that the internal energy vanishes, then the Hamiltonian in this model is linear, of the form (3)-(4) with $V_{int}(m)=0$.
+$(P+P_{\rm back})v=T.$ (33)
+Here, $P_{\rm back}$ is a background pressure that we treat as a physical parameter of the system.
+If the molecules do not interact so that the internal energy vanishes, then the Hamiltonian in this model is linear, of the form (3)-(4) with $V_{\rm int}(m)=0$.
 Then, a straightforward calculation following [15] shows that the equilibrium free energy $G$ in the ideal gas approximation is
-$G=-T\ln v=T\ln((P+P_{back})/T).$
+$G=-T\ln v=T\ln((P+P_{\rm back})/T).$
 We warn the reader that in this model the internal energy per particle vanishes, as opposed to the standard model of the ideal gas where it equals $\frac{3}{2}T/v$.
-Assume now that the temperature jumped from $T_{0}$ to $T_{1}>T_{0}$ and the background pressure simultaneously jumped from $P_{back}=0$ to $P_{back}=c>0$, while the volume is kept fixed.
+Assume now that the temperature jumped from $T_{0}$ to $T_{1}>T_{0}$ and the background pressure simultaneously jumped from $P_{\rm back}=0$ to $P_{\rm back}=c>0$, while the volume is kept fixed.
 One can think that the gas is contained in a chamber with one of the walls being a piston of area $a$, and one instantly changes the temperature and simultaneously switches on a force $f$ applied to the piston and directed towards the gas so that $c=f/a$.
-Given the parameters $T_{0}$, $T_{1}$ and $c$, there exist unique values of the volume $v$ and the initial pressure $P_{0}$ when such a jump corresponds to the ultrafast process, see Figure 1.
+Given the parameters $T_{0}$, $T_{1}$ and $c$, there exist unique values of the volume $v$ and the initial pressure $P_{0}$ when such a jump corresponds to the ultrafast process.
 This happens if, after the jump, the system lands in the equilibrium of the terminal system, so that
@@
 In particular, if the background pressure jump $c\ll 1$ is small, and $T_{1}-T_{0}=O(1)$, the volume $v$ is large in accordance to the ideal gas approximation.
-Figure 1: $T_{0}=1,T_{1}=5,c=2$
+Recall that this ultrafast process corresponds to the Reeb chord connecting initial and terminal equilibrium Legendrians.
+In the $(v,P)$-plain this chord corresponds to the intersection point of the Lagrangian projections of these Legendrians.
+This is visualized on Figure 1.
+Figure 1: Ultrafast process for $T_{0}=1,T_{1}=5,c=2$
 Note that the change in the Gibbs free energy is negative if the pressure jump $c<T_{1}-T_{0}$:
@@
 Remark 6.1.
+Our model ignores the Maxwell–Boltzmann distribution of particle velocities.
+A possible objection to this simplification is that the Maxwell–Boltzmann distribution depends on temperature, so a temperature jump would inevitably drive the system out of equilibrium.
+To address this, we tacitly assume that relaxation of the velocities to equilibrium is much faster than that of the coordinates.
+The physical feasibility of this assumption will be discussed elsewhere.
+Remark 6.2.
 When discussing an instantaneous isochoric process, as in the Stirling engine, temperature acts as a parameter, while pressure is a macroscopic variable.
-Introducing an artificial parameter, the background pressure, such that the total pressure is given by $P+P_{back}$, allows us to describe the simultaneous jump of two parameters: $T$ jumps from $T_{0}$ to $T_{1}$, and $P_{back}$ from 00 to $c$.
+Introducing an artificial parameter, the background pressure, such that the total pressure is given by $P+P_{\rm back}$, allows us to describe the simultaneous jump of two parameters: $T$ jumps from $T_{0}$ to $T_{1}$, and $P_{\rm back}$ from 0 to $c$.
 More generally, this trick applies to processes in which intensive variables exhibit instantaneous jumps while the corresponding extensive variables remain unchanged, see Section 6.2 for another example.
@@
 Let us explain now how such ultrafast process corresponds to a Reeb chord between two Legendrian submanifolds.
-We work in the reduced thermodynamic phase space $(z,p,q)$ where $z=-G,p=v,q=-P$.
+We work in the reduced thermodynamic phase space $(z,p,q)$ where z=G,p=v,q=Pz=-G,p=v,q=-P.
 This notation is in accordance with the fact that volume $v$ is an extensive variable and the pressure $P$ is the intensive variable [^5] We hope that the difference between $p$ and $P$ will not cause confusion for the reader..
 The Legendrian submanifold corresponding to the ideal gas is given by
-$\Lambda(T,P_{back})=\{(z,p,q)\in{\mathbb{R}}^{3}\;:\;z=\phi_{T}(q-P_{back}),\;p=\phi^{\prime}_{T}(q-P_{back})\}\;,$ (33)
+$\Lambda(T,P_{\rm back})=\{(z,p,q)\in{\mathbb{R}}^{3}\;:\;z=\phi_{T}(q-P_{\rm back}),\;p=\phi^{\prime}_{T}(q-P_{\rm back})\}\;,$ (37)
 where $\phi_{T}(q)=-T\ln(-q/T)$. Let us set $\Lambda_{0}=\Lambda(T_{0},0)$ and $\Lambda_{1}=\Lambda(T_{1},c)$.
@@
 Thus, the ultrafast process discussed above, corresponds exactly to the Reeb chord connecting $\Lambda_{0}$ with $\Lambda_{1}$.
-Furthermore, the orientation of the chord is given by $\partial/\partial z$ since $z_{1}-z_{0}>0$ by (32) and, as we recall, $z=-G$.
+Furthermore, the orientation of the chord is given by $\partial/\partial z$ since $z_{1}-z_{0}>0$ by (36) and, as we recall, z=Gz=-G.
 Let us illustrate the Reeb chord after the change of variables
@@
 $\psi(\overline{q})=\phi_{T_{1}}(\overline{q}-c)-\phi_{T_{0}}(\overline{q})\;.$
-The projections of the resulting Legendrians and the chord onto the $(q,z)$ plane (the front projection) are given on Fig. 3.
-Figure 3: $T_{0}=1,T_{1}=5,c=2$
+The projections of the resulting Legendrians and the chord onto the $(q,z)$ - plane (the front projection) are given on Fig. 3.
+Figure 3: Reeb chord for $T_{0}=1,T_{1}=5,c=2$
 Remark 6.3.
@@
 One can show that these circles are incomparable with respect to the partial order.
-This can be proved with the help of Legendrian spectral invariants which are monotone with respect to non-negative Legendrian isotopies, see [7, Section 4.1].
+This can be proved with the help of Legendrian spectral invariants which are monotone with respect to non-negative Legendrian isotopies, see [8, Section 4.1].
 Example 6.4.
 Here, we illustrate that Reeb chords can also correspond to a slow thermodynamic path, still on the ideal gas example.
-Suppose the temperature increases slowly with time $0\leq t\leq 1$, given by $T(t)=T_{0}+(T_{1}-T_{0})t$, and the external pressure also increases slowly $P_{ext}(t)=ct$, while the volume remains constant, $v(t)=v$.
+Suppose the temperature increases slowly with time $0\leq t\leq 1$, given by $T(t)=T_{0}+(T_{1}-T_{0})t$, and the external pressure also increases slowly $P_{\rm ext}(t)=ct$, while the volume remains constant, $v(t)=v$.
 The Lagrangian projection of the equilibrium Legendrian at time $t$ onto the $(p,q)$-plane is given by
@@
 In addition, the negative free energy increases during this process:
-$z(t)=-G(t)=T(t)\ln v(t)=(T+(T_{1}-T_{0})t)\ln v,\quad v\gg 1.$
+$z(t)=-G(t)=T(t)\ln v(t)=\bigl(T+(T_{1}-T_{0})t\bigr)\ln v,\quad v\gg 1.$
 Thus, the individual trajectory of the macroscopic state $(z_{0},p_{*},q_{*})$ is a Reeb chord.
@@
 Recall (cf.
-also Remark 6.1) that the total magnetic field is given by $H+H_{back}$, where $H_{back}$ is the background magnetic field, playing a role of the parameter.
-Suppose that the temperature of the Curie-Weiss magnet jumped from $T_{0}$ to $T_{1}>T_{0}$, and that the background magnetic field jumped from 00 to $c\in{\mathbb{R}}$.
-The reduced Legendrian equilibrium submanifold $\Lambda(T,H_{back})$ is defined by the first two equations in (26).
+also Remark 6.2) that the total magnetic field is given by $H+H_{\rm back}$, where $H_{\rm back}$ is the background magnetic field, playing a role of the parameter.
+Suppose that the temperature of the Curie-Weiss magnet jumped from $T_{0}$ to $T_{1}>T_{0}$, and that the background magnetic field jumped from 0 to $c\in{\mathbb{R}}$.
+The reduced Legendrian equilibrium submanifold $\Lambda(T,H_{\rm back})$ is defined by the first two equations in (26).
 Let us set $\Lambda_{0}=\Lambda(T_{0},0)$, and let $\Lambda_{1}=\Lambda(T_{1},c)$.
@@
 Proof.
-The Legendrian $\Lambda(T,H_{back})$ can be equivalently represented as
-$z=\phi_{T}(q+H_{back}+bp)-\frac{bp^{2}}{2},~{}~{}p=\phi_{T}^{\prime}(q+H_{back}+bp).$ (38)
+The Legendrian $\Lambda(T,H_{\rm back})$ can be equivalently represented as
+$z=\phi_{T}(q+H_{\rm back}+bp)-\frac{bp^{2}}{2},~~p=\phi_{T}^{\prime}(q+H_{\rm back}+bp).$ (42)
 Here, we have set
@@
 Let now $g_{t}$, $t\in[0,1]$ be a compactly supported contact isotopy of $J^{1}{\mathbb{R}}$ satisfying (41).
-Fix sufficiently large positive numbers $R>r>0$, let $A_{R}$ be the cube $[-R,R]^{3}\subset{\mathbb{R}}^{3}=J^{1}{\mathbb{R}}$ and denote by $S_{R}$ the circle ${\mathbb{R}}/R{\mathbb{Z}}$ in the $Q$-variable.
+Fix sufficiently large positive numbers $R>r>0$, let $A_{R}$ be the cube $R,R]^{3}\subset{\mathbb{R}}^{3}=J^{1}{\mathbb{R}}$ and denote by $S_{R}$ the circle ${\mathbb{R}}/R{\mathbb{Z}}$ in the $Q$-variable.
 For a Legendrian submanifold $Y$ which coincides with $K$ outside $A_{r}$, we denote by $\widehat{Y}$ the Legendrian submanifold of $J^{1}S_{R}$ obtained by the following procedure (see Fig.
@@
 inside the interior of $A_{R}$ without creating new chords from the zero section $L$ to the resulting submanifold $\widehat{Y}$.
-Then, glue together the points of this submanifold corresponding to $Q=-R$ and $Q=R$.
+Then, glue together the points of this submanifold corresponding to Q=RQ=-R and $Q=R$.
 This is possible simply because $L_{c}$ does not depend on $Q$.
@@
 ∎
+6.3 An ultrafast (de)magnetization for the Curie-Weiss model
+Consider the Curie-Weiss magnet at temperature $T$. Assume that the background magnetic field vanishes. Put $\psi(x)=T\ln\left(2\cosh(x/T)\right)$. Recall from Example 3.3 that in the equilibrium
+$z=\psi(q+bp)-\frac{bp^{2}}{2}\;,$ (47)
+$p=\psi^{\prime}(q+bp)\;,$ (48)
+and
+q=bp+Tu(p),q=-bp+Tu(p)\;, (49)
+with
+$u(p)=\frac{1}{2}\ln\frac{1+p}{1-p}\;.$
+We view $z$ and $p$ as functions of $b$, with $T$ and $q$ being fixed. Differentiating (47) by $b$ and substituting (48) we get that
+$\frac{dz}{db}=p^{2}/2\;.$
+Furthermore, by (49)
+$T^{-1}b=\frac{u(p)-T^{-1}q}{p}\;.$
+Differentiating by $p$ we get
+$T^{-1}\frac{db}{dp}=\frac{\left(pu^{\prime}(p)-u(p)\right)+T^{-1}q}{p^{2}}\;.$
+The expression in brackets in the nominator is the $y$-intercept of the function $u$ taken with the opposite sign.
+Since for $p>0$ the function $u$ is strictly convex with $u(0)=0$, its intercept is negative.
+It follows that for $p>0$, $dz/db>0$ and $db/dp>0$ (and hence $dp/db>0$).
+In particular, if as the result of certain ultrafast process, the exchange parameter (called sometimes coupling) $b$ is decreasing, the free energy increases and magnetization drops; and if $b$ is increasing, the free energy decreases and the magnetization is increasing.
+In some magnetic systems the manipulation of the parameter $b$ can be achieved by ultrashort laser pulses applied to the system ( see [29, 30]; we note that none of these papers use mean-field models; we introduce them here as an approximation.)
+Data availability statement. This paper does not use any data.
+Conflict of interest statement. On behalf of all authors, the corresponding author states that there is no conflict of interest.